### v1

# MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities

Abstract
Medical image segmentation has recently demonstrated impressive progress with deep neural networks, yet the heterogeneous modalities and scarcity of mask annotations limit the development of segmentation models on unannotated modalities. This paper investigates a new paradigm for leveraging generative models in medical applications: controllably synthesizing data for unannotated modalities, without requiring registered data pairs. Specifically, we make the following contributions in this paper: (i) we collect and curate a large-scale radiology image-text dataset, MedGen-1M, comprising modality labels, attributes, region, and organ information, along with a subset of organ mask annotations, to support research in controllable medical image generation; (ii) we propose a diffusion-based data engine, termed MRGen, which enables generation conditioned on text prompts and masks, synthesizing MR images for diverse modalities lacking mask annotations, to train segmentation models on unannotated modalities; (iii) we conduct extensive experiments across various modalities, illustrating that our data engine can effectively synthesize training samples and extend MRI segmentation towards unannotated modalities. The code, model, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
Figure 1: Motivations and Overview. Left: The heterogeneity of MRI modalities poses challenges to the generalization of segmentation models. Our proposed data engine, MRGen, addresses this by controllably synthesizing training data, enabling segmentation towards mask-unannotated modalities. Right: (a) Previous generative models rely on mask-annotated data and are restricted to data augmention in training modalities; (b) Image translation typically requires registered data pairs (indicated by dashed lines), limited to conversions between specific modalities; (c) MRGen introduces a new paradigm, enabling controllable generation across multiple modalities without relying on registered data pairs or mask-annotated data from target modalities. Different colors represent distinct modalities.
*: These authors contribute equally to this work.
$\dagger$: Corresponding author.
1 Introduction
Medical image segmentation [43, 21, 33, 55] has achieved remarkable success, becoming a cornerstone of intelligent healthcare. However, challenges such as data privacy, modality complexity, and the high cost of segmentation annotations significantly restrict its development towards unannotated modalities. As depicted in Figure 1 (left), these limitations are especially acute in MRI, which exhibits significant discrepancies across modalities.
Magnetic Resonance Imaging (MRI) is a non-invasive imaging technique widely recognized for its high-quality and radiation-free characteristics. However, MRI scans are often expensive, time-consuming, and sensitive to motion. Additionally, there is significant variability across different MRI modalities, and even within the same modality, differences in machines and scanning parameters can lead to substantial discrepancies in images. The complexity and variability of MRI, coupled with high annotation costs, lead to a scarcity of registered data pairs and segmentation annotations, posing significant challenges to the development of robust segmentation models. In this paper, we aim to leverage the progress in generative models, specifically, diffusion models [18], to facilitate the training of segmentation models on unannotated MRI modalities.
Generally speaking, there are three challenges when applying the generative models to MRI data synthesis: (i) paradigm: as presented in Figure 1 (right), previous models [13, 3] primarily focus on generating data for modalities with annotations. This approach limits their scalability to under-annotated or unannotated MRI modalities; (ii) data: the scarcity of MR images and fine-grained annotations makes it challenging to develop general and controllable generative models; (iii) controllability: to better benefit downstream tasks with synthetic data, generative models must enable controllable generation, guided by conditions such as text prompts or segmentation masks, to produce training samples for downstream tasks.
To tackle the above challenges, we collect and curate a large-scale radiology image-text dataset, MedGen-1M, featuring CT and MRI of various modalities sourced from the Internet and open-source data. Our dataset comprises 1.2 million 2D slices with modality labels, attributes, region, and organ information, with a subset also containing organ mask annotations. The extensive image-text pairs across diverse modalities provide a foundation for training a general MRI generation model, and the available mask annotations further enable more controllable generation.
Building on this foundation, this paper introduces MRGen, a diffusion-based data engine for controllable MRI synthesis, particularly, for modalities with no segmentation masks ever available. At training time, we employ a two-stage strategy: (i) text-guided pretraining on diverse, large-scale radiology image-text pair data, enabling the model to generate images across various modalities based on text descriptions; and (ii) mask-conditioned finetuning on data with mask annotations, facilitating controllable generation based on organ masks. Consequently, such a strategy even allows MRGen to extend its controllable generation abilities towards modalities that do not have segmentation annotations available. During inference, MRGen takes text prompts and organ masks as inputs, producing MR images aligned with the given modality, region, and organs. With such synthetic samples, we can thus facilitate MRI segmentation tools towards modalities without manual annotations.
Our contributions in this paper can be summarized as follows: (i) we establish a new paradigm for generative model applications in medical analysis, i.e., synthesizing data of unannotated modalities to serve as training samples for downstream segmentation models; (ii) we curate a large-scale radiology image-text dataset, MedGen-1M, which features detailed modality labels, attributes, regions, organ information, and a subset with organ mask annotations, providing a foundation for building general and controllable image generation models in the radiology domain; (iii) we develop MRGen, a diffusion-based data engine capable of generation conditioned on text prompts and masks, synthesizing MR images of various annotation-scarce modalities; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably synthesize high-quality MR images, boosting segmentation performance on unannotated modalities.
2 Related works
Generative models have been a research focus in computer vision for years, with GANs [12] and diffusion models [18, 45] leading the advancements. These models have found extensive applications across various tasks, including text-to-image generation [42, 24, 38, 48], image-to-image translation [4, 22, 58], artistic creation [30, 51, 44], and even challenging video generation [10, 19]. Notably, CycleGAN [58] employs cycle-consistency loss to facilitate image translation with unpaired data, while the Stable Diffusion series [42, 39, 11] efficiently produces high-resolution images in latent space, earning broad recognition.
Medical image synthesis aims to leverage generative models to improve medical image analysis models or tackle challenges such as data scarcity and privacy concerns [28]. Prior works train generative models on X-ray [3], CT [13], and brain MRI [8, 36], while methods like DiffTumor [5] and FreeTumor [49] focus specifically on generating tumor images to improve tumor segmentation.
Existing works [3, 13] primarily focus on in-domain augmentation within mask-annotated training modalities, and often struggle to generalize to unseen and unannotated modalities. To this end, this paper explores a novel paradigm for medical image generation, producing samples for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios.
Medical image segmentation has been a long-standing research topic, with various architectures being proposed [43, 21, 15, 56, 34]. Recently, inspired by SAM [27, 41], large-scale interactive segmentation models [33, 55, 9] have been developed for medical image analysis. However, the heterogeneity of MRI challenges the generalization of current segmentation models, which struggle with varying intensity distributions among diverse modalities. Existing methods generally address this problem by learning domain-invariant content, which either requires mixing training data from multiple domains, or relies on deliberately designed data augmentations strategies [57, 20, 46, 37, 50, 6]. Despite recent efforts in annotating segmentation masks for extensive MRI data [7, 14], building datasets with wide-ranging modalities, scanners and protocols is time-consuming and labor-intensive. In this paper, we explore controllable generative models to synthesize mask-conditioned MR images for any segmentation model, without tedious hyperparameter selection in data augmentation.
3 Data Curation
To train our proposed data engine, we first collect and curate a large-scale radiology dataset, termed MedGen-1M, containing CT and MR images with rich modality information, clinical attributes, and mask annotations. In the following, we elaborate on our data processing pipeline and present the statistics of the curated dataset.
Data collection. We first collect a large amount of abdominal CT and MR images from the Radiopaedia111radiopaeida.org website, licensed under CC BY-NC-SA 3.0222To obtain the data, readers must first get permission from Radiopaedia, then we will share the download link.. This portion of data comprises a wide variety of modalities, serving as the large-scale image-text pairs ($\{(\mathcal{I},\mathcal{T})\}$), with each sample containing an image and its modality label. As listed in Table 1, we also integrate CT and MR data with modality labels and abdominal organ mask annotations from several open-source datasets, to construct data triplets ($\{(\mathcal{I},\mathcal{T},\mathcal{M})\}$), which will be further explained in Section 4.
Automatic annotations. Given the wide variability in abdominal imaging, such as differences between the upper abdominal region and the pelvic region, relying solely on modality labels may not sufficiently distinguish these images. Thus, we divide the abdomen into six regions, including: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. We then use off-the-shelf BiomedCLIP [52] to categorize all 2D slices into these six categories to serve as region information.
Moreover, considering that even medical-specific text encoders may not clearly distinguish fine-grained correlations and differences between modalities like T1 and T2 [47, 53], we employ GPT-4 [1] to map all modality labels to free-text modality attributes that describe signal intensities of fat, muscle, and water. For example, T1 modality can be represented as fat high signal, muscle intermediate signal, water low signal. This aims to help the model understand the imaging characteristics among distinct modalities.
Discussion. After the data processing mentioned above, we assemble our MedGen-1M dataset. As shown in Table 1, the dataset consists of approximately 16K 3D volumes, covering CT and over a hundred MR modalities, totaling nearly 1.2M 2D slices. Each sample is paired with its corresponding modality labels, modality attributes, region, and organ information, with 190K samples featuring paired mask annotations. Compared to the existing datasets, the scale, modality diversity, and fine-grained annotations of MedGen-1M provide more information for training generative models tailored for medical imaging. More statistics will be detailed in Section B.2 of the Appendix.
Dataset Modality #Volumes #Slices #Masks
Radiopaedia-MR MR 5,414 205,039
Radiopaedia-CT CT 8,734 812,756
LiQA [31] MR 30 2,185 1,446
PanSeg [54] MR 766 33,360 13,779
PROMISE12 [29] MR 50 1,377 778
MSD-Prostate [2] MR 64 1,204 366
CHAOS-MRI [25] MR 60 1,917 1,492
MSD-Liver [2] CT 131 58,638 19,163
AMOS22CT [23] CT 300 41,430 30,279
Total CT & MR 15,499 1,156,529 186,525
Table 1: Dataset Statistics of MedGen-1M.
Figure 2: Architecture Overview. Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from $\mathcal{D}_{u}$, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on the mask-annotated source-domain dataset $\mathcal{D}_{s}$ and unannotated target-domain dataset $\mathcal{D}_{t}$, enabling controllable generation based on both text prompts and masks.
4 Method
We first formulate the problem in Section 4.1, followed by a detailed description of the proposed MRGen architecture in Section 4.2; then we elaborate on the training details of our model in Section 4.3; lastly, we present the procedure of synthesizing and filtering samples with our data engine for downstream segmentation tasks in Section 4.4.
4.1 Problem Formulation
In this paper, we focus on developing a diffusion-based controllable data engine for medical imaging, particularly for synthesizing data across various modalities with no mask annotations available. Specifically, our proposed MRGen ($\Phi_{\mathrm{MRGen}}$) generates the radiology image ($\mathcal{I}$), conditioned on a text prompt ($\mathcal{T}$) and the organs mask ($\mathcal{M}$), i.e., $\mathcal{I}=\Phi_{\mathrm{MRGen}}(\mathcal{T},\mathcal{M};\Theta,\Theta_{c})$, where $\Theta$ and $\Theta_{c}$ refer to the trainable parameters of the foundation generative model and the mask condition controller, respectively.
As described in Section 3, our large-scale dataset, i.e., $\mathcal{D}=\{\mathcal{D}_{u},\mathcal{D}_{s},\mathcal{D}_{t}\}$, encompasses three types of data, supporting the development of the proposed data engine. Concretely, the training process is carried out in two phases: (i) we pretrain a text-guided generative model on the image-text data ($\mathcal{D}_{u}=\{(\mathcal{I}_{u},\mathcal{T}_{u})\}$), covering diverse modalities; and (ii) we then train a condition controller jointly on image-text pairs with mask annotations from source-domain dataset ($\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$), and image-text pairs from the target-domain dataset ($\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$). The ultimate goal is to generate target-domain images ($\mathcal{I^{\prime}}_{t}$), conditioned on text prompts ($\mathcal{T^{\prime}}_{t}$) and masks ($\mathcal{M^{\prime}}_{t}$), to serve as training samples for segmentation models.
Relations to existing tasks. This paper makes two key advancements over existing methods: (i) conventional generative augmentation methods are confined to generating samples within mask-annotated training modalities [13, 3, 8]. Conversely, our model can synthesize data for modalities lacking mask annotations, conditioned on text prompts and masks, improving segmentation models on these challenging scenarios; (ii) unlike image translation works [58, 22, 36] that typically require registered data pairs for training or are limited to specific conversions, our model provides flexible and controllable generation with no need for registered data or mask-annotated data of the target modality.
4.2 Architecture
For our generation paradigm, we expect the model to leverage the large number of image-text data pairs, and the limited data with segmentation masks, to achieve controllable generation for segmentation-scarce modalities. Our proposed model (MRGen) consists of three components: (i) latent encoding; (ii) text-guided generation; and (iii) mask-conditioned generation, as detailed below.
Latent encoding. Given the high resolution of medical images, we adopt the idea of latent diffusion [42], for efficient generation within a compressed, low-dimensional latent space. As shown in Figure 2 (a), we utilize an autoencoder architecture, where a 2D MRI slice ($\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$) is encoded into a latent code ($\mathbf{z}\in\mathbb{R}^{h\times w\times d}$) for diffusion training. The decoder then reconstructs the image ($\hat{\mathcal{I}}$) from the latent code, as expressed by the following:
^$\hat{\mathcal{I}}=\phi_{\mathrm{Dec}}(\mathbf{z})=\phi_{\mathrm{Dec}}(\phi_{\mathrm{Enc}}(\mathcal{I}))$
Text-guided generation. This part follows the diffusion model paradigm, consisting of a forward diffusion process and a denoising process. Concretely, the forward process progressively adds noise to the latent features ($\mathbf{z}_{0}$) over $T$ steps towards white Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,1)$. At any intermediate timestep $t\in[1,T]$, the noisy features ($\mathbf{z}_{t}$) can be expressed as: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ and $\bar{\alpha}_{t}$ represent predefined hyperparameters, which will be detailed in Section A of the Appendix.
As depicted in Figure 2 (b), the learnable denoising process typically reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$. To model this process in the compressed latent space, we adopt a UNet [43]. Specifically, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”. We employ an off-the-shelf text encoder ($\phi_{\mathrm{text}}$) to encode the text prompt into embeddings, expressed as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$. These text embeddings are then fed into our model via cross-attention, serving as the key and value, with visual latent codes as the query.
Mask-conditioned generation. Next, we incorporate mask conditions to enable more controllable generation. As shown in Figure 2 (c), we adopt the idea similar to ControlNet [51], i.e., initializing the mask encoder ($\phi_{\mathrm{mask}}$) with the weights of the diffusion UNet encoder, coupled with a learnable downsampling module ($\phi_{\mathrm{down}}$) to align the dimensions of the mask with the latent code. In the input mask ($\mathcal{M}\in\mathbb{R}^{H\times W\times 1}$), we use distinct intensity values to indicate the different organs, and encode it as a residual into the corresponding layers of the UNet decoder, guiding the generation process. In addition, zero convolution initialization is employed to stabilize the training process, as observed in [51]. For each layer $\mathcal{F}$ of the diffusion UNet decoder, the output $\mathbf{O}$ can now be formulated as follows:
$\mathbf{O}=\mathcal{F}(\mathbf{z}_{t})+\phi_{\mathrm{mask}}(\mathbf{z}_{t},\phi_{\mathrm{down}}(\mathcal{M}),\phi_{\mathrm{text}}(\mathcal{T}))$
4.3 Model Training
Here, we present the training procedure for our generative model, including (i) autoencoder reconstruction, (ii) text-guided pretraining, and (iii) mask-conditioned finetuning.
Autoencoder reconstruction. The autoencoder for compression is trained solely on raw images from $\mathcal{D}_{u}$, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to a VAE [26] and $\gamma$ represents a predefined weight.
Text-guided pretraining. The diffusion-based generative model, parameterized by $\Theta$, is trained on a large number of image-text pairs in $\mathcal{D}_{u}$, covering diverse modalities. The objective function can be formulated as the MSE loss between the added Gaussian noise ($\boldsymbol{\epsilon}$) and the prediction ($\hat{\boldsymbol{\epsilon}}$):
$\mathcal{L}=\mathbb{E}_{\mathcal{D}_{u},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}(\mathbf{z}_{t},t,\mathcal{T})||_{2}^{2}\Big{]}$
This pretraining phase enables MRGen to generate radiology images across diverse modalities based on text prompts.
Mask-conditioned finetuning. The mask condition controller, consisting of the mask encoder ($\phi_{\mathrm{mask}}$) and the learnable downsampling module ($\phi_{\mathrm{down}}$), can be trained jointly on the source-domain triplet dataset, $\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$, and target-domain dataset comprising only image-text pairs, $\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$, with all other parameters frozen. The training objective $\mathcal{L}_{c}$ can be expressed as:
$\mathcal{L}_{c}=\mathbb{E}_{\mathcal{D}_{s},\mathcal{D}_{t},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{c}(\mathbf{z}_{t},t,\mathcal{T},\mathcal{M})||_{2}^{2}\Big{]}$
Here, incorporating target-domain data without mask annotations prevents overfitting to the source domain.
Discussion. (i) The large-scale image-text dataset ($\mathcal{D}_{u}$) comprises a wide range of modalities, covering those of source-domain dataset ($\mathcal{D}_{s}$) and target-domain dataset ($\mathcal{D}_{t}$); (ii) Mask-conditioned finetuning on mask-annotated data allows MRGen to learn correlations between images and organ masks; (iii) Such a two-stage training strategy then enables MRGen to extend its controllable generation ability towards target domains lacking mask annotations.
Figure 3: Synthetic Data Construction Pipeline. MRGen takes text prompt and mask as conditions for controllably generating radiology images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
4.4 Synthetic Data for Segmentation Training
With our data engine, we aim to generate target-domain data conditioned on the segmentation masks, to facilitate training segmentation models towards unannotated modalities.
Data synethesis. At inference time, we feed the text prompt ($\mathcal{T^{\prime}}_{t}$) and organ mask ($\mathcal{M^{\prime}}_{t}$) as conditions into our MRGen model ($\Phi_{\mathrm{MRGen}}$) to generate the corresponding target-domain image ($\mathcal{I^{\prime}}_{t}$). For simplicity, when constructing the input conditions, we directly utilize region, organ information, and mask annotations from source-domain data.
Data autofilter. To ensure the fidelity of generated images to the mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2 [41] model, as depicted in Figure 3. Specifically, we feed the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) into SAM2 to predict a segmentation confidence score ($s_{\mathrm{conf}}$), and a pseudo mask annotation used to calculate the IoU score ($s_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}_{t}$. A sample is considered high-quality and aligned with the mask condition if both its IoU score and confidence score exceed the predefined thresholds; otherwise, it is discarded.
Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [57] CycleGAN [58] MRGen (Ours)
CM. T1 CM. T2-SPIR 156.98 228.16 157.77 44.82
CM. T2-SPIR CM. T1 156.98 261.97 188.91 60.79
MP. T2 MP. ADC 305.56 422.73 112.82 99.38
MP. ADC MP. T2 305.56 416.31 190.82 123.55
PS. T1 PS. T2 76.95 120.36 237.52 34.35
PS. T2 PS. T1 76.95 126.27 89.26 58.90
LQ. T1 CM. T2-SPIR 109.46 281.56 182.62 88.76
CM. T2-SPIR LQ. T1 109.46 246.73 260.06 106.45
MP. ADC PR. T2 387.29 434.54 221.27 116.35
PR. T2 MP. ADC 387.29 365.10 140.72 88.43
AM. CT CM. T1 148.04 226.23 248.06 42.53
AM. CT CM. T2-SPIR 164.73 259.96 268.49 65.30
ML. CT CM. T2-SPIR 176.51 265.64 307.79 96.72
Average FID $\downarrow$ 197.06 281.20 200.47 78.95
Table 2: Quantitative Results (FID) on Generation. Smaller values indicate better performance. Here, CM., MP., PS., LQ., PR., AM., and ML., denote CHAOS-MRI, MSD-Prostate, PanSeg, LiQA, PROMISE12, AMOS22, and MSD-Liver, respectively.
Source Dataset Source Modality Target Dataset Target Modality DualNorm [57] nnUNet [21] UMamba [34]
$\mathcal{D}_{s}$ CycleGAN MRGen $\mathcal{D}_{s}$ CycleGAN MRGen Oracle Box SAM2 nnUNet [21] $\mathcal{D}_{t}$
Same Datasets, Different Modalities
CHAOS-MRI T1 CHAOS-MRI T2-SPIR 14.00 6.90 7.58 66.18 4.02 10.58 67.35 45.45 53.12 83.90
CHAOS-MRI T2-SPIR CHAOS-MRI T1 12.50 0.80 1.38 58.10 0.62 0.22 57.24 43.48 51.94 90.60
MSD-Prostate T2 MSD-Prostate ADC 1.43 5.52 40.92 57.83 0.19 45.06 52.58 61.50 65.39 82.35
MSD-Prostate ADC MSD-Prostate T2 12.94 22.20 57.06 61.95 11.80 62.00 64.05 61.07 66.40 89.80
PanSeg T1 PanSeg T2 0.21 0.68 2.40 9.78 0.38 2.13 9.34 31.69 41.11 82.28
PanSeg T2 PanSeg T1 0.11 0.30 3.59 12.07 0.27 5.08 10.29 30.94 41.18 80.27
Different Datasets, Different Modalities
LiQA T1 CHAOS-MRI T2-SPIR 19.23 16.20 7.84 31.73 11.05 8.65 37.30 62.72 70.86 70.83
CHAOS-MRI T2-SPIR LiQA T1 31.09 15.80 41.02 57.28 10.33 41.22 52.54 56.11 67.99 93.60
MSD-Prostate ADC PROMISE12 T2 1.43 17.19 42.13 35.33 23.71 43.24 37.12 50.41 53.61 87.00
PROMISE12 T2 MSD-Prostate ADC 9.84 19.20 59.95 56.88 21.75 57.21 49.77 61.50 65.39 82.35
Different Datasets, Different Modalities, CT - MRI
AMOS22 CT CHAOS-MRI T1 19.78 0.11 6.75 22.50 0.05 8.06 26.73 43.48 51.94 90.60
AMOS22 CT CHAOS-MRI T2-SPIR 16.09 8.88 52.49 56.23 3.19 43.21 60.53 45.45 53.12 83.90
MSD-Liver CT CHAOS-MRI T2-SPIR 1.58 3.12 10.14 38.67 1.65 11.06 40.93 62.72 70.86 70.83
Average DSC score 10.79 8.99 25.63 43.43 6.85 25.89 43.52 50.50 57.92 83.72
Table 3: Quantitative Results (DSC score) on Segmentation. Oracle Box denotes the DSC score compared between the perturbed oracle box prompts for SAM2, and the ground truth; $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$ denote training with real and manually annotated source-domain and target-domain data respectively. Results are organized by three different cross-modality simulations, where the best results under nnUNet and UMamba are bolded respectively, while the second best are underlined.
5 Experiments
Here, we start by describing the experimental settings in Section 5.1; then, in Section 5.2 and 5.3, we present a comprehensive evaluation of our method from both quantitative and qualitative perspectives; lastly, we provide results for ablation experiments in Section 5.4 to prove the effectiveness of our proposed methods.
5.1 Experimental Settings
We aim to evaluate our proposed data engine from two aspects: (i) image generation, and (ii) segmentation. Concretely, to simulate MRI segmentation towards unannotated modalities, we construct 8 cross-modal dataset pairs within our MedGen-1M. Each pair comprises a mask-annotated source-domain dataset and an unannotated target-domain dataset. The generative model is trained on each dataset pair to synthesize target-domain samples, which are then used to train segmentation models. We assess the quality of generated images, and the performance of the segmentation models on the target-domain test set.
Baselines. For image generation, we compare mask-conditioned generated images from MRGen against two other approaches: CycleGAN [58] for translating source-domain images to the target domain; and DualNorm [57] for aggressive augmentation of source-domain images. For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) MRGen generated data; and (v) manually annotated target-domain data, serving as a strong oracle baseline. We adopt nnUNet [21] and UMamba [34] as segmentation frameworks for all settings, except for DualNorm, which uses a customized UNet. Moreover, we include SAM2 [41] as a reference for interactive semi-automatic segmentation, using randomly perturbed oracle boxes as prompts.
Evaluation metrics. We employ different metrics for assessment: For image generation, we adopt the commonly used Fréchet Inception Distance score (FID) [16] to assess the diversity and quality of the generated images. For the downstream segmentation models, we employ the widely used Dice Similarity Coefficient (DSC) [35] to compare the predicted segmentation mask and ground truth. Considering segmentation consistency, we stack slices back into 3D volumes for DSC calculation.
Figure 4: Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, prostate, liver, right kidney, left kidney and spleen are contoured with different colors.
Implementation details. All the images are resized to resolution 512 $\times$ 512 pixels. For training, we start by training the VAE with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations. Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively. Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance for better performance. All training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [32] optimizer. The downsampling factor, latent feature dimension $d$, KL divergence loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively. During inference, we perform 50-step sampling using DDIM [45], with the classifier-free guidance [17] weight set to $w=7.0$. We adopt the pretrained text encoder in BiomedCLIP [52] to encode text prompts, and SAM2 [41] for automatic filtering. For each conditional mask, we generate 20 candidates and select the best two that satisfy the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively.
Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. We visualize prostate on the middle row, and liver, right kidney, left kidney and spleen on the upper and lower row with different colors.
5.2 Quantitative Results
Image generation. As shown in Table 2, images from the source domain exhibit high FID values with the target domain, indicating substantial discrepancies between distinct radiology modalities. Similarly, images augmented by DualNorm and translated by CycleGAN also exhibit high FID values, confirming their limited ability to emulate target-domain images. Conversely, our MRGen presents a significantly lower FID, demonstrating its capability to accurately generate images in target modalities, thus providing a foundation for training segmentation models with synthetic data.
Segmentation. As presented in Table 3, we have the following four key observations: (i) The significant discrepancies among different modalities hinder the ability of nnUNet and UMamba trained exclusively on source-domain data to generalize to the target domain, resulting in the lowest average DSC scores. (ii) The conventional image translation approach, CycleGAN, and the augmentation-based method, DualNorm, achieve moderate improvements in average DSC scores. However, these methods still struggle in the target domain and occasionally underperform in specific scenarios. (iii) In contrast, our proposed MRGen generates high-quality images for target domains, achieving the highest DSC score in 11 out of 13 experiments and significantly outperforming CycleGAN. Notably, MRGen consistently improves performance across all settings compared to source-only training, demonstrating the adaptability and effectiveness of its synthetic data across various segmentation architectures. (iv) The interactive SAM2 model, trained on substantially larger datasets, exhibits robust zero-shot segmentation capabilities with perturbed oracle box prompts. However, as a semi-automatic method, its reliance on spatial prompts and manual intervention limits its scalability and practical applicability.
5.3 Qualitative Results
Generation. As depicted in Figure 4, the images of distinct modalities exhibit substantial visual discrepancies, making it challenging for DualNorm to simulate via augmentation. Additionally, CycleGAN is prone to instability during training, frequently leading to model collapse when learning such complex translations. Conversely, our proposed MRGen effectively generates images that closely resemble the target domain and align with conditioned organ masks. This capability provides a solid foundation for the controlled generation of training data to extend segmentation models towards unannotated modalities.
Segmentation. As shown in Figure 5, despite significant variability between the source and target domains, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data. However, the data derived from DualNorm and CycleGAN lead to unsatisfactory segmentation results, due to notable discrepancies from the target domain or low quality.
Method AutoFilter $\mathcal{D}_{t}$ Image CHAOS-MRI [25] MSD-Prostate [2]
T1 $\rightarrow$ T2S. T2S. $\rightarrow$ T1 T2 $\rightarrow$ ADC ADC $\rightarrow$ T2
nnUNet [21] 6.90 0.80 5.52 22.20
nnUNet (MRGen) 16.53 15.10 39.90 18.92
22.30 20.27 42.79 25.34
30.16 29.01 49.04 40.89
66.18 58.10 57.83 61.95
Table 4: Ablation Study on Segmentation Performance (DSC score). Here, T2S. denotes T2-SPIR.
5.4 Ablation Studies
To validate the effectiveness of our modules and strategies, we conduct ablation studies on the downstream segmentation task. Specifically, we train nnUNet using data synthesized under different training and inference settings. As presented in Table 4, we have the following observations: (i) even without integrating unannotated target-domain data during training, MRGen still improves segmentation performance on the target modality. This indicates that the two-stage training strategy effectively enables MRGen to extend controllable generation to modalities without mask annotations; (ii) the autofilter pipeline based on SAM2 further improves performance by selecting data pairs that align better with the mask conditions. More ablation studies on image generation will be provided in Section D.2 of the Appendix.
6 Conclusion
This paper explores a novel paradigm of applying generative models in medical imaging: controllably synthesizing data for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios. To support this, we curate a large-scale radiology image-text dataset, MedGen-1M, featuring detailed modality labels, attributes, regions, and organ information, along with a subset of organ mask annotations. Built upon the dataset, our diffusion-based controllable data engine, MRGen, generates MR images conditioned on text prompts and masks, enabling data synthesis across unannotated modalities for training segmentation models. Comprehensive quantitative and qualitative evaluations across diverse modalities demonstrate that MRGen can effectively synthesize data, advancing MRI segmentation for unannotated modalities.
Appendix
Appendix A Preliminaries on Diffusion Models
Diffusion Models [18] are a class of deep generative models that convert Gaussian noise into structured data samples through an iterative denoising process. These models typically comprise a forward diffusion process and a reverse denoising process.
Specifically, the forward diffusion process progressively introduces Gaussian noise into an image ($\mathbf{x}_{0}$) via a Markov process over $T$ steps. Let $\mathbf{x}_{t}$ represent the noisy image at step $t$. The transition from $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t}$ can be formulated as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$
Here, $\beta_{t}\in(0,1)$ represents pre-determined hyperparameters that control the variance at each step. By defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, the properties of Gaussian distributions and the reparameterization trick allow for a refined expression:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$
This insight provides a concise expression for the forward process with Gaussian noise $\boldsymbol{\epsilon}$ as: $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$.
Diffusion models also encompass a reverse denoising process that reconstructs images from noise. A UNet-based model [43] is typically utilized to learn the reverse diffusion process $p_{\theta}$, represented as:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$
Here, $\mu_{\theta}$ represents the predicted mean of Gaussian distribution, derived from the estimated noise $\boldsymbol{\epsilon}_{\theta}$ as:
$\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t))$
Building on this foundation, Latent Diffusion Models [42] adopt a Variational Autoencoder (VAE [26]) to project images into a learned, compressed, low-dimensional latent space. The forward diffusion and reverse denoising processes are then performed on the latent codes ($\mathbf{z}$) within this latent space, significantly reducing computational cost and improving efficiency.
Appendix B Details of MedGen-1M
In this section, we provide additional details about our collected and curated MedGen-1M dataset. In Section B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Section B.2, we present more comprehensive statistics about the dataset.
B.1 Automatic Annotations
We employ an automated annotation pipeline to annotate the data in our MedGen-1M, ensuring that the samples contain comprehensive clinically relevant annotation information. This process primarily includes two components: human body region classification and modality explanation, which will be detailed as follows.
Region classification. Considering the wide range and variability of abdominal imaging, we adopt the off-the-shelf BiomedCLIP [52] image encoder to encode all 2D slices and the BiomedCLIP text encoder to encode predefined text descriptions of six abdominal regions. Based on the cosine similarity between the image and text embeddings, the 2D slices are classified into these six categories, including Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. For text encoding, we use a templated text prompt as input:
{mdframed}
[backgroundcolor=gray!10, linewidth=0pt] This is a radiology image that shows \$region\$ of a human body, and probably contains \$organ\$.
Here, \$region\$ and \$organ\$ represent the items in the following list:
{mdframed}
[backgroundcolor=gray!10, linewidth=0pt] (region, organ) = [ (‘Upper Thoracic Region’, ‘lung, ribs and clavicles’), (‘Middle Thoracic Region’, ‘lung, ribs and heart’), (‘Lower Thoracic Region’, ‘lung, ribs and liver’), (‘Upper Abdominal Region’, ‘liver, spleen, pancreas, kidney and stomach’), (‘Lower Abdominal Region’, ‘kidney, small intestine, colon, cecum and appendix’), (‘Pelvic Region’, ‘rectum, bladder, prostate/uterus and pelvic bones’) ]
Modality explanation. To capture the correlations and distinctions among various modality labels, we leverage GPT-4 [1] to generate free-text descriptions detailing the signal intensities of fat, muscle, and water for each modality label. This helps the model better understand the imaging characteristics of distinct modalities. The prompt we use is as follows:
{mdframed}
[backgroundcolor=gray!10, linewidth=0pt] As a senior doctor and medical imaging researcher, please help me map radiological imaging modalities to the signal intensities of fat, muscle, and water, as well as their corresponding brightness levels. Please provide the answer in the following format: fat {} signal, muscle {} signal, water {} signal, fat {}, muscle {}, water {}. Now, tell me the attributes of \$modality\$.
B.2 Dataset Statistics
In this section, we present more detailed statistics about our curated MedGen-1M dataset, including the unannotated image-text pairs from Radiopaedia333radiopaeida.org, as well as the mask-annotated data sourced from various open-source datasets.
Data without mask annotations. For the image-text pairs from Radiopaedia-CT and Radiopaedia-MRI, which are used for training the autoencoder and text-guided generation, we allocate 1% of the data as a test set to evaluate reconstruction and generation performance, maximizing the amount of data available for pretraining. As a result, 1,007,616 samples are used for training, comprising 804,628 from Radiopaedia-CT and 202,988 from Radiopaedia-MRI. The test set consists of 10,179 samples, with 8,128 from Radiopaedia-CT and 2,051 from Radiopaedia-MRI.
We conduct a statistical analysis of the distribution of modalities in Radiopaedia-MRI, as presented in Figure 6 (a). The free-text modality labels cover approximately 300 categories, providing a diverse set of MRI modalities that form a crucial foundation for MRGen to learn text-guided generation and expand its mask-conditioned generation capabilities towards modalities originally lacking mask annotations. Furthermore, the distribution of images across different regions in Radiopaedia-CT and Radiopaedia-MRI is presented in Figure 6 (b).
Figure 6: Data Statistics of Radiopaedia. (a) Distribution of slice counts across various modalities in Radiopaedia-MRI; (b) Proportional distribution of slices across different regions in Radiopaedia-CT and Radiopaedia-MRI.
Data with mask annotations. Following the SAT [55], we split the data with mask annotations into training and test sets, as detailed in Table 5. For dataset pairs comprising different datasets, we use their shared organs as the segmentation targets.
Dataset Organs Modality Train Test
# Vol. # Slc. # Slc. w/ mask # Vol. # Slc. # Slc. w/ mask
LiQA [31] Liver T1-weighted 24 1,718 1,148 6 467 298
PanSeg [54] Pancreas T1-weighted 309 14,656 5,961 75 3,428 1,400
T2-weighted 305 12,294 5,106 77 2,982 1,312
MSD-Prostate [2] Prostate T2-weighted 26 492 100 6 110 83
ADC 26 492 100 6 110 83
PROMISE12 [29] Prostate T2-weighed 40 1,137 645 10 110 133
CHAOS-MRI [25] Liver, Right Kidney, T1-weighted 32 1,018 770 8 276 230
Left Kidney, Spleen T2-SPIR 16 503 388 4 120 104
MSD-Liver [2] Liver CT 105 46,695 15,260 26 11943 3903
AMOS22CT [23] Liver, Right Kidney, CT 200 26,069 19,023 100 15,361 11,256
Left Kidney, Spleen, etc.*
Total / / 1,083 105,074 48,501 218 34,907 18,802
Table 5: Details of Segmentation-annotated Datasets in MedGen-1M. Here, # Vol. represents the number of 3D Volumes, # Slc. denotes the number of 2D slices, and # Slc. w/ mask indicates the number of 2D slices with mask annotations. *: AMOS22CT [23] also contains images and mask annotations for the following organs: Gallbladder, Esophagus, Stomach, Aorta, Inferior Vena Cava, Pancreas, Duodenum, Urinary Bladder, and Adrenal Gland.
Appendix C Implementation Details
In this section, we will provide a comprehensive explanation of the implementation details discussed in the paper. Specifically, Section C.1 describes the preprocessing and augmentation strategies applied to the training data. Section C.2 elaborates on the details of the autofilter pipeline. Finally, Section C.3 outlines the implementation details of the baselines.
C.1 Preprocessing & Augmentation
Data preprocessing. To ensure consistency across data from various sources and modalities, we apply tailored preprocessing strategies as follows: (i) For data from Radiopaedia, the images are directly rescaled to the range [0, 1]; (ii) For CT images with mask annotations, intensities are clipped to [-300, 200] and then rescaled to [0, 1]; (iii) For MR images with mask annotations, intensities are clipped to the 0.5 and 99.5 percentiles and rescaled to [0, 1]. After normalization, all data are subsequently rescaled to [-1, 1] for training various components of MRGen, including the autoencoder, diffusion UNet, and mask condition controller. For training downstream segmentation models, images are rescaled to [0, 255] and saved in ‘.png’ format, followed by the preprocessing configurations of nnUNet [21] and UMamba [34].
Data augmentation. During autoencoder training, we apply random data augmentations to the images with a 20% probability. These augmentations included horizontal flipping, vertical flipping, and rotations of 90, 180, 270. In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller. For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [21] and UMamba [34].
C.2 Autofilter Pipeline
When deploying our proposed data engine, MRGen, to synthesize data for segmentation model training, we adopt the off-the-shelf SAM2 [41] model to perform automatic interactive segmentation on generated images, with the mask conditions as prompt. This pipeline automatically selects samples that are fidelity to the condition masks, thus ensuring the quality of synthetic image-mask pairs. In this section, we elaborate on more implementation details of this automatic filtering pipeline, particularly focusing on the generation of MR images that encompass multiple organs.
Specifically, we begin by defining the following thresholds: confidence threshold ($\tau_{\mathrm{conf}}$), IoU score threshold ($\tau_{\mathrm{IoU}}$), average confidence threshold ($\bar{\tau}_{\mathrm{conf}}$), and average IoU threshold ($\bar{\tau}_{\mathrm{IoU}}$). Both the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) are fed into SAM2. For each organ mask $\mathcal{M^{\prime}}^{i}_{t}$ in $\mathcal{M^{\prime}}_{t}$, SAM2 will output a confidence score ($s^{i}_{\mathrm{conf}}$) and a pseudo mask annotation, which is then used to calculate the IoU score ($s^{i}_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}^{i}_{t}$. For each generated sample ($\mathcal{I^{\prime}}_{t}$), it is regarded as a high-quality sample if the following conditions are satisfied: $\{s^{i}_{\mathrm{IoU}}\geq\tau_{\mathrm{IoU}},s^{i}_{\mathrm{conf}}\geq\tau_{\mathrm{conf}}\,|\,\forall i\}$, and $\{\bar{s}_{\mathrm{IoU}}\geq\bar{\tau}_{\mathrm{IoU}}$, $\bar{s}_{\mathrm{conf}}\geq\bar{\tau}_{\mathrm{conf}}\}$. Otherwise, the sample will be discarded.
In all experiments, the predefined thresholds are set as follows: $\tau_{\mathrm{IoU}}=0.70$, $\tau_{\mathrm{conf}}=0.80$, $\bar{\tau}_{\mathrm{IoU}}=0.80$, and $\bar{\tau}_{\mathrm{conf}}=0.90$.
C.3 Baselines
CycleGAN. We follow the official implementation and training strategies across all experimental settings. Subsequently, source-domain images are translated into the target domain and paired with the source-domain masks to create paired samples for training downstream segmentation models.
SAM2. We use SAM2-Large in all the experiments. Segmentation results are derived in a slice-by-slice and organ-by-organ manner: For each slice with mask annotations, we simulate box prompts for each annotated organ individually. To take into account the error introduced by manual intervention, the oracle boxes are randomly shifted at each corner, by up to 8% of the image resolution, following the protocol in MedSAM [33].
DualNorm. Following the official implementation, we apply random non-linear augmentation on each image in the source domain, to generate a source-dissimilar training sample, and train the dual-normalization model. All preprocessing steps, network architectures, and training strategies follow the official recommendations, with the exception that images are rescaled to 512 $\times$ 512, consistent with other approaches. Note that we evaluate DualNorm on all the slices in the test set, offering a more rigorous evaluation compared to the official implementation, which only considers slices with segmentation annotation.
Appendix D More Experiments
In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine. In Section D.1, we showcase both quantitative and qualitative results of in-domain generation. In Section D.2, we conduct extensive ablation studies to illustrate the improvements introduced by our design in radiology image generation. Finally, in Section D.3, we provide additional qualitative results to validate the accuracy and flexibility of the generated outputs.
D.1 In-domain Generation
Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domain. Moreover, as presented in Table 6, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated source-domain data. This offers a feasible solution to address concerns about medical data privacy.
Dataset Source Modality Target Modality $\mathcal{D}_{s}$ $\mathcal{D}_{t}$
$\mathcal{D}_{s}$ MRGen $\mathcal{D}_{s}$ MRGen $\mathcal{D}_{t}$
CHAOS-MRI [25] T1 T2-SPIR 90.60 88.14 4.02 67.35 83.90
T2-SPIR T1 83.90 82.06 0.62 57.24 90.60
Table 6: More Quantitative Results (DSC score) on Segmentation. We compare the performance of nnUNet [21] trained on real data versus synthetic data generated by our MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
Moreover, we provide visualizations of in-domain generation in Figure 7, qualitatively demonstrating that our MRGen can reliably perform controllable generation of a large number of samples within the training domain with mask annotations.
Figure 7: Qualitative Results of In-domain Generation.
D.2 Ablation Studies
In this section, we provide more ablation study results on image generation, to demonstrate the necessity of model and data design within our proposed data engine, focusing on autoencoder reconstruction and text-guided generation.
Autoencoder reconstruction. We evaluate the reconstruction performance of various VAEs on the test set derived from our curated MedGen-1M dataset, comprising Radiopaedia-CT and Radiopaedia-MRI. The evaluated models include: (i) the VAE of pretrained Stable Diffusion (VAE-SDM), (ii) the VAE of Stable Diffusion finetuned on our dataset (Finetuned-VAE-SDM), (iii) our model trained solely on Radiopaedia-MRI data (VAE-MRGen (MRI-only)), and (iv) our VAE-MRGen trained on the complete dataset (VAE-MRGen (Ours)).
The reconstruction quality is assessed using standard metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), L1 Error, and Mean Squared Error (MSE) between the reconstructed and original images.
As illustrated in Table 8, we draw the following three key observations: (i) Finetuning on radiology data: VAEs finetuned on radiology data significantly improve the reconstruction quality compared to pretrained Stable Diffusion models, revealing the limited generalizability of natural image VAEs to radiology images. (ii) Latent code dimension: our autoencoders with a larger latent code feature dimension achieve further improvements in reconstruction quality, suggesting that increasing the increased capacity of latent code enables better representations. (iii) Extra data: incorporating CT data into training further improves reconstruction quality, approaching the performance of lossless compression. To this end, unlike RoentGen [3] that relies on pretrained VAEs, we employ high-capacity autoencoders trained specifically on radiology data, to ensure that the generative models built upon this foundation can produce better image details.
Text-guided generation. We evaluate the text-guided image generation performance of various models on the test set derived from our MedGen-1M dataset, encompassing Radiopaedia-CT and Radiopaedia-MRI. The evaluated models include: (i) the pretrained Stable Diffusion (SDM), (ii) the Stable Diffusion finetuned on our MedGen-1M (Finetuned-SDM), (iii) our model using only free-text modality labels as text conditions (MRGen (Modality-only)), and (iv) our MRGen, which utilizes templated text prompts incorporating modality labels, attributes, regions, and organs information (MRGen (Ours)).
We employ commonly used evaluation metrics for image generation, including the Fréchet Inception Distance (FID [16]) to assess the quality and diversity of generated images. Additionally, we utilize BiomedCLIP [52] to calculate two similarity metrics: image-to-image similarity (CLIP-I) between the generated images and ground truth, and image-to-text similarity (CLIP-T) between the generated images and the free-text modality descriptions.
As presented in Table 8, we draw the following three observations: (i) Finetuning on radiology data: pretrained Stable Diffusion tends to generate natural images rather than ideal radiology images, resulting in suboptimal performance across all evaluation metrics. However, finetuning on our dataset significantly enhances its generative performance. (ii) Text encoder: using BiomedCLIP [52] as the text encoder, trained on biomedical data, instead of CLIP [40] in Stable Diffusion, improves MRGen’s ability to capture similarities and differences among radiology modalities. Additionally, our high-capacity VAE, provides further potential improvements to generation quality. (iii) Templated text prompt: incorporating templated text prompts with detailed information (including modality, attributes, regions, and organs) outperforms using modality labels alone. This enables MRGen to better grasp the relationships among distinct modalities, thereby further improving generation quality, particularly in terms of image-to-image similarity with the ground truth.
Model Dim PSNR $\uparrow$ SSIM $\uparrow$ L1 Error $\downarrow$ MSE $\downarrow$
VAE-SDM [42] 4 31.32 0.989 0.0294 0.0037
Finetuned-VAE-SDM 4 35.65 0.996 0.0190 0.0014
VAE-MRGen (MRI-only) 16 40.56 0.998 0.0115 0.0005
VAE-MRGen (Ours) 16 42.62 0.999 0.0091 0.0003
Table 7: Ablation Studies on Autoencoder Reconstruction.
Model FID $\downarrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SDM [42] 249.24 0.3151 0.1748
Finetuned-SDM 91.48 0.6698 0.3199
MRGen (Modality-only) 41.82 0.7512 0.3765
MRGen (Ours) 39.63 0.8457 0.3777
Table 8: Ablation Studies on Text-guided Generation.
D.3 More Qualitative Results
In this section, we provide qualitative visualizations of more datasets, covering both image generation and segmentation.
Image generation. As depicted in Figure 9 and 10, we present additional visualizations of controllable generation on target modalities lacking mask annotations. These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
Image segmentation. As presented in Figure 11 and 12, we provide more visualizations of segmentation models trained using synthetic data on modalities that originally lack mask annotations. This validates that the samples generated by MRGen can effectively assist in training segmentation models, achieving impressive performance in previously unannotated scenarios.
Appendix E Limitations & Future Works
E.1 Limitations
Our proposed data engine, MRGen, is not without its limitations. Specifically, MRGen encounters difficulties when generating conditioned on extremely small organ masks and occasionally produces false-negative samples.
Extremely small organ masks. The morphology of the same organ, such as the liver or spleen, can vary significantly across different slices of a 3D volume, resulting in significant variability in their corresponding masks. Furthermore, the distribution of these masks is often imbalanced, with extremely small masks being relatively rare. When generating in the latent space, these masks are further downsampled, leading to unstable generation quality, as depicted in Figure 8 (a). A feasible solution to mitigate this issue is to increase the amount of data with mask annotations, thereby improving the model’s robustness.
False-negative samples. Another challenge arises from the varying number of organs to be segmented on each slice. For instance, one slice may contain the liver, kidneys, and spleen, while another may include only the liver and spleen. This variability causes MRGen to occasionally generate additional segmentation targets not specified in the mask condition. For example, as illustrated in Figure 8 (b), kidneys are unexpectedly synthesized by MRGen, despite not being included in the mask conditions, leading to false negatives during the training of downstream segmentation networks. A feasible solution is to design a more comprehensive and robust data filtering pipeline to filter these false-negative samples. Alternatively, simple manual selection can serve as a quick and effective method to remove samples that do not meet the requirements.
Figure 8: Failure Cases Analysis. Our proposed MRGen is not without limitations: (a) it may struggle to handle extremely small organ masks; (b) it occasionally produces false-negative samples, such as the unexpected synthesis of kidneys in the given example.
E.2 Future Works
To address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated CT data or collecting more MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
Figure 9: More Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 10: More Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 11: More Qualitative Results on Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 12: More Qualitative Results on Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.

### v2

@@
-MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities
+MRGen: Segmentation Data Engine For Underrepresented MRI Modalities
 Abstract
-Medical image segmentation has recently demonstrated impressive progress with deep neural networks, yet the heterogeneous modalities and scarcity of mask annotations limit the development of segmentation models on unannotated modalities.
-This paper investigates a new paradigm for leveraging generative models in medical applications: controllably synthesizing data for unannotated modalities, without requiring registered data pairs.
-Specifically, we make the following contributions in this paper: (i) we collect and curate a large-scale radiology image-text dataset, MedGen-1M, comprising modality labels, attributes, region, and organ information, along with a subset of organ mask annotations, to support research in controllable medical image generation; (ii) we propose a diffusion-based data engine, termed MRGen, which enables generation conditioned on text prompts and masks, synthesizing MR images for diverse modalities lacking mask annotations, to train segmentation models on unannotated modalities; (iii) we conduct extensive experiments across various modalities, illustrating that our data engine can effectively synthesize training samples and extend MRI segmentation towards unannotated modalities.
-The code, model, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
+Training medical image segmentation models for rare yet clinically significant imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire.
+This paper investigates leveraging generative models to synthesize training data, to train segmentation models for underrepresented modalities, particularly on annotation-scarce MRI.
+Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset having pixelwise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks.
+MRGen can generate realistic images for diverse MRI modalities lacking mask annotations, facilitating segmentation training in low-source domains; (iii) extensive experiments across multiple modalities demonstrate that MRGen significantly improves segmentation performance on unannotated modalities by providing high-quality synthetic data.
+We believe that our method bridges a critical gap in medical image analysis, extending segmentation capabilities to scenarios that are challenging to acquire manual annotations.
+The codes, models, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
 Figure 1: Motivations and Overview.
-Left: The heterogeneity of MRI modalities poses challenges to the generalization of segmentation models.
-Our proposed data engine, MRGen, addresses this by controllably synthesizing training data, enabling segmentation towards mask-unannotated modalities.
-Right: (a) Previous generative models rely on mask-annotated data and are restricted to data augmention in training modalities; (b) Image translation typically requires registered data pairs (indicated by dashed lines), limited to conversions between specific modalities; (c) MRGen introduces a new paradigm, enabling controllable generation across multiple modalities without relying on registered data pairs or mask-annotated data from target modalities.
-Different colors represent distinct modalities.
+Left: The heterogeneity of MRI modalities challenges the generalization of segmentation models.
+Our proposed data engine, MRGen, overcomes this by controllably synthesizing training data for segmentation models.
+Right: (a) Prior generative models are restricted to data augmentation for well-annotated modalities; (b) Image translation typically requires registered data pairs (dashed lines), and is limited to specific modality conversions; (c) MRGen enables controllable generation across diverse modalities, creating data for training segmentation models towards underrepresented modalities.
+Distinct colors represent different modalities.
 *: These authors contribute equally to this work.
@@
 1 Introduction
-Medical image segmentation [43, 21, 33, 55] has achieved remarkable success, becoming a cornerstone of intelligent healthcare.
-However, challenges such as data privacy, modality complexity, and the high cost of segmentation annotations significantly restrict its development towards unannotated modalities.
-As depicted in Figure 1 (left), these limitations are especially acute in MRI, which exhibits significant discrepancies across modalities.
-Magnetic Resonance Imaging (MRI) is a non-invasive imaging technique widely recognized for its high-quality and radiation-free characteristics.
-However, MRI scans are often expensive, time-consuming, and sensitive to motion.
-Additionally, there is significant variability across different MRI modalities, and even within the same modality, differences in machines and scanning parameters can lead to substantial discrepancies in images.
-The complexity and variability of MRI, coupled with high annotation costs, lead to a scarcity of registered data pairs and segmentation annotations, posing significant challenges to the development of robust segmentation models.
-In this paper, we aim to leverage the progress in generative models, specifically, diffusion models [18], to facilitate the training of segmentation models on unannotated MRI modalities.
-Generally speaking, there are three challenges when applying the generative models to MRI data synthesis: (i) paradigm: as presented in Figure 1 (right), previous models [13, 3] primarily focus on generating data for modalities with annotations.
-This approach limits their scalability to under-annotated or unannotated MRI modalities; (ii) data: the scarcity of MR images and fine-grained annotations makes it challenging to develop general and controllable generative models; (iii) controllability: to better benefit downstream tasks with synthetic data, generative models must enable controllable generation, guided by conditions such as text prompts or segmentation masks, to produce training samples for downstream tasks.
-To tackle the above challenges, we collect and curate a large-scale radiology image-text dataset, MedGen-1M, featuring CT and MRI of various modalities sourced from the Internet and open-source data.
-Our dataset comprises 1.2 million 2D slices with modality labels, attributes, region, and organ information, with a subset also containing organ mask annotations.
-The extensive image-text pairs across diverse modalities provide a foundation for training a general MRI generation model, and the available mask annotations further enable more controllable generation.
-Building on this foundation, this paper introduces MRGen, a diffusion-based data engine for controllable MRI synthesis, particularly, for modalities with no segmentation masks ever available.
-At training time, we employ a two-stage strategy: (i) text-guided pretraining on diverse, large-scale radiology image-text pair data, enabling the model to generate images across various modalities based on text descriptions; and (ii) mask-conditioned finetuning on data with mask annotations, facilitating controllable generation based on organ masks.
-Consequently, such a strategy even allows MRGen to extend its controllable generation abilities towards modalities that do not have segmentation annotations available.
-During inference, MRGen takes text prompts and organ masks as inputs, producing MR images aligned with the given modality, region, and organs.
-With such synthetic samples, we can thus facilitate MRI segmentation tools towards modalities without manual annotations.
-Our contributions in this paper can be summarized as follows: (i) we establish a new paradigm for generative model applications in medical analysis, i.e., synthesizing data of unannotated modalities to serve as training samples for downstream segmentation models; (ii) we curate a large-scale radiology image-text dataset, MedGen-1M, which features detailed modality labels, attributes, regions, organ information, and a subset with organ mask annotations, providing a foundation for building general and controllable image generation models in the radiology domain; (iii) we develop MRGen, a diffusion-based data engine capable of generation conditioned on text prompts and masks, synthesizing MR images of various annotation-scarce modalities; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably synthesize high-quality MR images, boosting segmentation performance on unannotated modalities.
+Medical image segmentation [50, 23, 39, 5] has shown remarkable success by training on extensive manual annotations, becoming a cornerstone of intelligent healthcare systems.
+However, developing models for underrepresented imaging modalities remains challenging, due to data privacy, modality complexity, and high cost of manual mask annotations [8, 16], especially for rare yet clinically important modalities, for example, Magnetic Resonance Imaging (MRI).
+Despite being non-invasive and radiation-free, MRI scanning is expensive and exhibits substantial variability across modalities and scanning protocols [43].
+This lack of standardization and numerous hyperparameters fragment the already limited dataset, challenging the development of robust segmentation models, as illustrated in Figure 1 (left).
+In this paper, we investigate the potential of generative models, particularly diffusion models, to synthesize MRI data for training segmentation models on underrepresented modalities.
+While generative models offer a promising solution, they face unique challenges: (i) data availability remains a significant obstacle.
+Existing approaches have primarily focused on data augmentation for well-annotated modalities such as X-ray [3] and CT [14, 15], as depicted in Figure 1 (right).
+However, MRI data is relatively scarce, and highly diverse across modalities, making it less explored for generative modeling; (ii) controllability is critical to facilitate downstream tasks such as segmentation.
+Thus, generative models must enable controllable synthesis based on conditions such as texts, masks, or both.
+Yet, prior works [55, 9, 42, 60, 32] cannot simultaneously support both conditions, limiting their abilitiy to control the generated modalities, regions, and organs effectively.
+Our first contribution is the collection of a large-scale, high-quality radiology image-text dataset, MRGen-DB (short for “Database for MRI Generation”), which includes MRI scans across various modalities sourced from the Internet and open-source repositories.
+The dataset consists of nearly 250,000 2D slices, enriched with detailed annotations such as modality labels, attributes, region, and organ information, with a subset providing organ masks.
+This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training a general MRI generation model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
+For controllable data generation, we present MRGen, a diffusion-based data engine for MRI synthesis, that supports conditioning on both text prompts and segmentation masks.
+We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) mask-conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks.
+Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enables training segmentation models for these underrepresented modalities with synthetic data.
+Overall, our contributions can be summarized as follows: (i) we explore the use of generative models for MRI synthesis across annotation-scarce modalities, facilitating training segmentation models for underrepresented modalities; (ii) we curate MRGen-DB, a large-scale radiology image-text dataset, which features detailed modality labels, attributes, regions, and organs information, with a subset of organ mask annotations, providing a robust foundation for medical generative modeling; (iii) we develop MRGen, a diffusion-based data engine capable of controllable generation, conditioned on templated text prompts and segmentation masks; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably generate high-quality MR images, improving ‘zero-shot’ segmentation performance for unannotated modalities.
+To our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first general medical generative model tailored for annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
 2 Related works
@@
 These models have found extensive applications across various tasks, including text-to-image generation [49, 25, 45, 57], image-to-image translation [4, 24, 69], artistic creation [35, 61, 51], and even challenging video generation [11, 21].
-Notably, CycleGAN [58] employs cycle-consistency loss to facilitate image translation with unpaired data, while the Stable Diffusion series [42, 39, 11] efficiently produces high-resolution images in latent space, earning broad recognition.
-Medical image synthesis aims to leverage generative models to improve medical image analysis models or tackle challenges such as data scarcity and privacy concerns [28].
-Prior works train generative models on X-ray [3], CT [13], and brain MRI [8, 36], while methods like DiffTumor [5] and FreeTumor [49] focus specifically on generating tumor images to improve tumor segmentation.
-Existing works [3, 13] primarily focus on in-domain augmentation within mask-annotated training modalities, and often struggle to generalize to unseen and unannotated modalities.
-To this end, this paper explores a novel paradigm for medical image generation, producing samples for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios.
-Medical image segmentation has been a long-standing research topic, with various architectures being proposed [43, 21, 15, 56, 34].
-Recently, inspired by SAM [27, 41], large-scale interactive segmentation models [33, 55, 9] have been developed for medical image analysis.
-However, the heterogeneity of MRI challenges the generalization of current segmentation models, which struggle with varying intensity distributions among diverse modalities.
-Existing methods generally address this problem by learning domain-invariant content, which either requires mixing training data from multiple domains, or relies on deliberately designed data augmentations strategies [57, 20, 46, 37, 50, 6].
-Despite recent efforts in annotating segmentation masks for extensive MRI data [7, 14], building datasets with wide-ranging modalities, scanners and protocols is time-consuming and labor-intensive.
-In this paper, we explore controllable generative models to synthesize mask-conditioned MR images for any segmentation model, without tedious hyperparameter selection in data augmentation.
-3 Data Curation
-To train our proposed data engine, we first collect and curate a large-scale radiology dataset, termed MedGen-1M, containing CT and MR images with rich modality information, clinical attributes, and mask annotations.
-In the following, we elaborate on our data processing pipeline and present the statistics of the curated dataset.
+Notably, CycleGAN [69] employs cycle-consistency loss to facilitate image translation with unpaired data, while Stable Diffusion series [49, 47, 12] efficiently produces high-resolution images in latent space, earning broad recognition.
+Medical image synthesis aims to leverage generative models to tackle challenges such as data scarcity [37, 32], biases [33], and privacy concerns [31].
+Prior works primarily focus on X-ray [3], CT [14, 15], and brain MRI [9, 42, 64], with approaches like DiffTumor [6] and FreeTumor [58] specifically targeting tumor generation to boost tumor segmentation.
+While these methods have proven effectiveness in data augmentation within well-annotated training modalities and regions, they still struggle to generalize to modalities lacking manual mask annotations.
+To this end, this paper investigates adopting generative models to facilitate more robust segmentation models with high-quality synthetic data of annotation-scarce modalities.
+Medical image segmentation has been a long-standing research topic, with various architectures proposed [50, 23, 17, 67, 40].
+Recently, inspired by SAM [30, 48], large-scale segmentation models [39, 66, 10] have been developed.
+However, the heterogeneity of MRI challenges the generalization of existing models, which struggle with intensity variations among diverse modalities.
+Existing methods attempt to address this with image translation [52, 46, 28] or relying on deliberately designed augmentation strategies to learn domain-invariant content [68, 22, 54, 44, 59, 7].
+In this paper, we explore controllable generative models to synthesize data for segmentation training, particularly towards underrepresented modalities lacking mask annotations, thus resembling a ‘zero-shot’ segmentation scenario.
+Figure 2: Architecture Overview.
+Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
+3 Method
+Here, we start to formulate the problem of interest in Sec.
+3.1, followed by a detailed description of the dataset curation procedure in Sec.
+3.2.
+Later, we elaborate on the proposed MRGen architecture and the training details of our model in Sec.
+3.3 and Sec.
+3.4, respectively.
+Lastly, we present the procedure of synthesizing and filtering samples for downstream segmentation tasks in Sec.
+3.5.
+3.1 Problem Formulation
+Given a text prompt ($\mathcal{T}$) describing modality, region, and organs, along with the organs mask ($\mathcal{M}$), our proposed MRGen ($\Phi_{\mathrm{MRGen}}$) enables to generate the MR image ($\mathcal{I}$):
+$\mathcal{I}=\Phi_{\mathrm{MRGen}}(\mathcal{T},\mathcal{M};\Theta,\Theta_{c})$
+where $\Theta$ and $\Theta_{c}$ refer to parameters of the generative model and the mask condition controller, respectively.
+Developing such a controllable data engine, thus enables synthesizing high-quality data to train segmentation models for the challenging ‘zero-shot’ scenario.
+Relations to existing tasks.
+Numerous studies have proven the effectiveness of generative models for data augmentation [55] on well-annotated modalities and regions, such as CT [15, 14], X-ray [3], and brain MRI [9, 42], however, this is not the focus of our work.
+Instead, we target the more challenging scenario of synthesizing MR images for scarce and underrepresented modalities where no manual mask annotations are available.
+While image translation methods [69, 27, 52, 46] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data for training or are limited to specific modality conversions.
+In contrast, our proposed MRGen framework offers a flexible and controllable generation pipeline, enabling the synthesis of complex abdominal MRI data across diverse modalities, even in the absence of mask annotations.
+Dataset # Volumes # Slices # Masks
+Radiopaedia-MRI 5,414 205,039
+PanSeg [65] 766 33,360 13,779
+MSD-Prostate [2] 64 1,204 366
+CHAOS-MRI [26] 60 1,917 1,492
+PROMISE12 [34] 50 1,377 778
+LiQA [36] 30 2,185 1,446
+Total 6,384 245,082 17,861
+Table 1: Dataset Statistics of MRGen-DB.
+3.2 Dataset Curation
+The scarcity of MR images with comprehensive text descriptions and mask annotations poses challenges for training generative models.
+To tackle this limitation, we present MRGen-DB (short for “Database for MRI Generation”), a meticulously curated large-scale radiology dataset featuring diverse MR images enriched with modality information, detailed clinical attributes, and precise mask annotations.
+Below, we detail our data processing pipeline and provide comprehensive dataset statistics.
 Data collection.
-We first collect a large amount of abdominal CT and MR images from the Radiopaedia111radiopaeida.org website, licensed under CC BY-NC-SA 3.0222To obtain the data, readers must first get permission from Radiopaedia, then we will share the download link..
-This portion of data comprises a wide variety of modalities, serving as the large-scale image-text pairs ($\{(\mathcal{I},\mathcal{T})\}$), with each sample containing an image and its modality label.
-As listed in Table 1, we also integrate CT and MR data with modality labels and abdominal organ mask annotations from several open-source datasets, to construct data triplets ($\{(\mathcal{I},\mathcal{T},\mathcal{M})\}$), which will be further explained in Section 4.
+The primary source of our dataset is abdominal MR images obtained from Radiopaedia [^1] Radiopaedia.org., licensed under CC BY-NC-SA 3.0222Access to this dataset requires prior permission from Radiopaedia, after which we will provide the download link..
+This portion of data includes a diverse array of imaging modalities, forming extensive image-text pairs suitable for training text-guided generative models.
+Each sample consists of an MR image and its corresponding free-text modality label.
+To enhance the dataset’s coverage and utility, we also augment it with abdominal MRI data from multiple open-source repositories.
+These additional data sources include modality labels and organ-specific mask annotations, forming comprehensive data triplets ($\{\mathcal{I},\mathcal{T},\mathcal{M}\}$).
+This augmentation enables more sophisticated controllable generation guided by both textual descriptions and anatomical masks, broadening the dataset’s applicability.
 Automatic annotations.
-Given the wide variability in abdominal imaging, such as differences between the upper abdominal region and the pelvic region, relying solely on modality labels may not sufficiently distinguish these images.
-Thus, we divide the abdomen into six regions, including: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region.
-We then use off-the-shelf BiomedCLIP [52] to categorize all 2D slices into these six categories to serve as region information.
-Moreover, considering that even medical-specific text encoders may not clearly distinguish fine-grained correlations and differences between modalities like T1 and T2 [47, 53], we employ GPT-4 [1] to map all modality labels to free-text modality attributes that describe signal intensities of fat, muscle, and water.
-For example, T1 modality can be represented as fat high signal, muscle intermediate signal, water low signal.
-This aims to help the model understand the imaging characteristics among distinct modalities.
+Abdominal imaging is highly variable, exhibiting significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region.
+Relying solely on modality labels is insufficient to differentiate these distinctions.
+To address this, we divide the abdomen into six anatomical regions: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region.
+Using the pre-trained BiomedCLIP model [62], we automatically categorize all 2D slices into these regions.
+To maintain annotation quality, slices with low confidence scores ($<40\%$) are intentionally left unlabeled.
+This process enriches the dataset with detailed region-specific information.
+Distinguishing fine-grained modality differences, such as between T1 and T2, presents additional challenges, even for advanced medical-specific text encoders [56, 63].
+To overcome this, we employ GPT-4 [1] to map modality labels into free-text attributes that describe the signal intensities of specific tissues, including fat, muscle, and water.
+For example, the T1 modality can be represented as: fat high signal, muscle intermediate signal, water low signal.
+These detailed descriptions enable the model to understand and differentiate imaging characteristics across modalities.
+To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a subset of the data.
+Specifically, 2% of the region annotations and 20% of the modality attributes have been reviewed, achieving high accuracies of 95.33% and 91.67%, respectively.
+This verification step ensures the high quality of the dataset and strengthens its applicability for downstream tasks.
 Discussion.
-After the data processing mentioned above, we assemble our MedGen-1M dataset.
-As shown in Table 1, the dataset consists of approximately 16K 3D volumes, covering CT and over a hundred MR modalities, totaling nearly 1.2M 2D slices.
-Each sample is paired with its corresponding modality labels, modality attributes, region, and organ information, with 190K samples featuring paired mask annotations.
-Compared to the existing datasets, the scale, modality diversity, and fine-grained annotations of MedGen-1M provide more information for training generative models tailored for medical imaging.
-More statistics will be detailed in Section B.2 of the Appendix.
-Dataset Modality #Volumes #Slices #Masks
-Radiopaedia-MR MR 5,414 205,039
-Radiopaedia-CT CT 8,734 812,756
-LiQA [31] MR 30 2,185 1,446
-PanSeg [54] MR 766 33,360 13,779
-PROMISE12 [29] MR 50 1,377 778
-MSD-Prostate [2] MR 64 1,204 366
-CHAOS-MRI [25] MR 60 1,917 1,492
-MSD-Liver [2] CT 131 58,638 19,163
-AMOS22CT [23] CT 300 41,430 30,279
-Total CT & MR 15,499 1,156,529 186,525
-Table 1: Dataset Statistics of MedGen-1M.
-Figure 2: Architecture Overview.
-Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from $\mathcal{D}_{u}$, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on the mask-annotated source-domain dataset $\mathcal{D}_{s}$ and unannotated target-domain dataset $\mathcal{D}_{t}$, enabling controllable generation based on both text prompts and masks.
-4 Method
-We first formulate the problem in Section 4.1, followed by a detailed description of the proposed MRGen architecture in Section 4.2; then we elaborate on the training details of our model in Section 4.3; lastly, we present the procedure of synthesizing and filtering samples with our data engine for downstream segmentation tasks in Section 4.4.
-4.1 Problem Formulation
-In this paper, we focus on developing a diffusion-based controllable data engine for medical imaging, particularly for synthesizing data across various modalities with no mask annotations available.
-Specifically, our proposed MRGen ($\Phi_{\mathrm{MRGen}}$) generates the radiology image ($\mathcal{I}$), conditioned on a text prompt ($\mathcal{T}$) and the organs mask ($\mathcal{M}$), i.e., $\mathcal{I}=\Phi_{\mathrm{MRGen}}(\mathcal{T},\mathcal{M};\Theta,\Theta_{c})$, where $\Theta$ and $\Theta_{c}$ refer to the trainable parameters of the foundation generative model and the mask condition controller, respectively.
-As described in Section 3, our large-scale dataset, i.e., $\mathcal{D}=\{\mathcal{D}_{u},\mathcal{D}_{s},\mathcal{D}_{t}\}$, encompasses three types of data, supporting the development of the proposed data engine.
-Concretely, the training process is carried out in two phases: (i) we pretrain a text-guided generative model on the image-text data ($\mathcal{D}_{u}=\{(\mathcal{I}_{u},\mathcal{T}_{u})\}$), covering diverse modalities; and (ii) we then train a condition controller jointly on image-text pairs with mask annotations from source-domain dataset ($\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$), and image-text pairs from the target-domain dataset ($\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$).
-The ultimate goal is to generate target-domain images ($\mathcal{I^{\prime}}_{t}$), conditioned on text prompts ($\mathcal{T^{\prime}}_{t}$) and masks ($\mathcal{M^{\prime}}_{t}$), to serve as training samples for segmentation models.
-Relations to existing tasks.
-This paper makes two key advancements over existing methods: (i) conventional generative augmentation methods are confined to generating samples within mask-annotated training modalities [13, 3, 8].
-Conversely, our model can synthesize data for modalities lacking mask annotations, conditioned on text prompts and masks, improving segmentation models on these challenging scenarios; (ii) unlike image translation works [58, 22, 36] that typically require registered data pairs for training or are limited to specific conversions, our model provides flexible and controllable generation with no need for registered data or mask-annotated data of the target modality.
+After completing the data processing above, we assemble the MRGen-DB dataset, which includes approximately 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as presented in Table 1.
+Each sample is paired with its modality label, attributes, region, and organ information, with about 18,000 samples featuring mask annotations.
+The scale, diversity, and fine-grained annotations of MRGen-DB provide sufficient information for training generative models tailored for MR images.
+More statistics are provided in Sec.
+B.2.
 3.3 Architecture
-For our generation paradigm, we expect the model to leverage the large number of image-text data pairs, and the limited data with segmentation masks, to achieve controllable generation for segmentation-scarce modalities.
-Our proposed model (MRGen) consists of three components: (i) latent encoding; (ii) text-guided generation; and (iii) mask-conditioned generation, as detailed below.
+Our research focus expects the model to leverage abundant image-text pairs and limited mask-annotated data to achieve controllable generation for underrepresented modalities.
+Specifically, our model (MRGen) comprises three components: (i) latent encoding; (ii) text-guided generation; and (iii) mask-conditioned generation, as detailed below.
 Latent encoding.
-Given the high resolution of medical images, we adopt the idea of latent diffusion [42], for efficient generation within a compressed, low-dimensional latent space.
-As shown in Figure 2 (a), we utilize an autoencoder architecture, where a 2D MRI slice ($\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$) is encoded into a latent code ($\mathbf{z}\in\mathbb{R}^{h\times w\times d}$) for diffusion training.
-The decoder then reconstructs the image ($\hat{\mathcal{I}}$) from the latent code, as expressed by the following:
+To handle high-resolution medical images, we first map them into a low-dimensional latent space for efficient training.
+As shown in Figure 2 (a), we employ an autoencoder, that encodes a 2D slice ($\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$) into a latent representation ($\mathbf{z}\in\mathbb{R}^{h\times w\times d}$), which can be reconstructed to image ($\hat{\mathcal{I}}$) by the decoder, expressed as:
 ^$\hat{\mathcal{I}}=\phi_{\mathrm{Dec}}(\mathbf{z})=\phi_{\mathrm{Dec}}(\phi_{\mathrm{Enc}}(\mathcal{I}))$
+To enable the generative model to effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
 Text-guided generation.
-This part follows the diffusion model paradigm, consisting of a forward diffusion process and a denoising process.
+This part follows the diffusion model paradigm, comprising a forward diffusion process and a denoising process.
 Concretely, the forward process progressively adds noise to the latent features ($\mathbf{z}_{0}$) over $T$ steps towards white Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,1)$.
-At any intermediate timestep $t\in[1,T]$, the noisy features ($\mathbf{z}_{t}$) can be expressed as: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ and $\bar{\alpha}_{t}$ represent predefined hyperparameters, which will be detailed in Section A of the Appendix.
-As depicted in Figure 2 (b), the learnable denoising process typically reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$.
-To model this process in the compressed latent space, we adopt a UNet [43].
-Specifically, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”.
-We employ an off-the-shelf text encoder ($\phi_{\mathrm{text}}$) to encode the text prompt into embeddings, expressed as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$.
-These text embeddings are then fed into our model via cross-attention, serving as the key and value, with visual latent codes as the query.
+At any intermediate timestep $t\in[1,T]$, the noisy visual features ($\mathbf{z}_{t}$) is expressed as: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$, and $\bar{\alpha}_{t}$ denotes predefined hyperparameters.
+As depicted in Figure 2 (b), the denoising process adopts a UNet [50] and reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$.
+Concretely, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”.
+These templated prompts ensure sufficient clinical information to distinguish distinct modalities, regions, and organs.
+We employ an off-the-shelf BiomedCLIP [62] text encoder ($\phi_{\mathrm{text}}$) to encode them into embeddings, denoted as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$.
+These embeddings are integrated into our model via cross-attention, serving as the key and value, with visual features ($\mathbf{z}_{t}$) as the query.
+The output ($\mathbf{O}_{\mathrm{cross}}$) of each cross-attention layer ($\mathcal{F}_{\mathrm{cross}}$) is represented as:
+$\mathbf{O}_{\mathrm{cross}}=\mathcal{F}_{\mathrm{cross}}(\mathbf{z}_{t},\phi_{\mathrm{text}}(\mathcal{T}))$
 Mask-conditioned generation.
-Next, we incorporate mask conditions to enable more controllable generation.
-As shown in Figure 2 (c), we adopt the idea similar to ControlNet [51], i.e., initializing the mask encoder ($\phi_{\mathrm{mask}}$) with the weights of the diffusion UNet encoder, coupled with a learnable downsampling module ($\phi_{\mathrm{down}}$) to align the dimensions of the mask with the latent code.
-In the input mask ($\mathcal{M}\in\mathbb{R}^{H\times W\times 1}$), we use distinct intensity values to indicate the different organs, and encode it as a residual into the corresponding layers of the UNet decoder, guiding the generation process.
-In addition, zero convolution initialization is employed to stabilize the training process, as observed in [51].
-For each layer $\mathcal{F}$ of the diffusion UNet decoder, the output $\mathbf{O}$ can now be formulated as follows:
-$\mathbf{O}=\mathcal{F}(\mathbf{z}_{t})+\phi_{\mathrm{mask}}(\mathbf{z}_{t},\phi_{\mathrm{down}}(\mathcal{M}),\phi_{\mathrm{text}}(\mathcal{T}))$
+We then incorporate mask conditions to enable more controllable generation.
+As presented in Figure 2 (c), we initialize the mask encoder ($\phi_{\mathrm{mask}}$) using weights from the encoder of the diffusion UNet pre-trained in the previous stage, coupled with a learnable downsampling module ($\phi_{\mathrm{down}}$) to align dimensions.
+The input mask ($\mathcal{M}\in\mathbb{R}^{H\times W\times 1}$) uses distinct intensity values to represent different organs, and is integrated as a residual into the UNet decoder.
+For each block ($\phi^{i}_{\mathrm{mask}}$) of the mask encoder, the output ($\mathbf{O}^{i}$) of the corresponding block ($\mathcal{F}^{i}$) in the diffusion UNet decoder, is formulated as:
+$\mathbf{O}^{i}=\mathcal{F}^{i}(\mathbf{z}_{t})+\phi_{\mathrm{mask}}^{i}(\mathbf{z}_{t},\phi_{\mathrm{down}}(\mathcal{M}),\phi_{\mathrm{text}}(\mathcal{T}))$
 3.4 Model Training
-Here, we present the training procedure for our generative model, including (i) autoencoder reconstruction, (ii) text-guided pretraining, and (iii) mask-conditioned finetuning.
+Here, we present the training procedure for our proposed model, including: (i) autoencoder reconstruction, (ii) text-guided pretraining, and (iii) mask-conditioned finetuning.
 Autoencoder reconstruction.
-The autoencoder for compression is trained solely on raw images from $\mathcal{D}_{u}$, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to a VAE [26] and $\gamma$ represents a predefined weight.
+The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [29] and $\gamma$ denotes a predefined weight.
 Text-guided pretraining.
-The diffusion-based generative model, parameterized by $\Theta$, is trained on a large number of image-text pairs in $\mathcal{D}_{u}$, covering diverse modalities.
-The objective function can be formulated as the MSE loss between the added Gaussian noise ($\boldsymbol{\epsilon}$) and the prediction ($\hat{\boldsymbol{\epsilon}}$):
-$\mathcal{L}=\mathbb{E}_{\mathcal{D}_{u},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}(\mathbf{z}_{t},t,\mathcal{T})||_{2}^{2}\Big{]}$
-This pretraining phase enables MRGen to generate radiology images across diverse modalities based on text prompts.
+The diffusion-based generative model, parameterized by $\Theta$, is trained on a large number of image-text pairs, covering diverse modalities.
+The objective function is formulated as the MSE loss between the added Gaussian noise ($\boldsymbol{\epsilon}$) and the prediction ($\hat{\boldsymbol{\epsilon}}$):
+$\mathcal{L}=\mathbb{E}_{t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}(\mathbf{z}_{t},t,\mathcal{T})||_{2}^{2}\Big{]}$
+This pretraining phase enables MRGen to generate MR images across various modalities based on text prompts.
 Mask-conditioned finetuning.
-The mask condition controller, consisting of the mask encoder ($\phi_{\mathrm{mask}}$) and the learnable downsampling module ($\phi_{\mathrm{down}}$), can be trained jointly on the source-domain triplet dataset, $\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$, and target-domain dataset comprising only image-text pairs, $\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$, with all other parameters frozen.
-The training objective $\mathcal{L}_{c}$ can be expressed as:
-$\mathcal{L}_{c}=\mathbb{E}_{\mathcal{D}_{s},\mathcal{D}_{t},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{c}(\mathbf{z}_{t},t,\mathcal{T},\mathcal{M})||_{2}^{2}\Big{]}$
-Here, incorporating target-domain data without mask annotations prevents overfitting to the source domain.
+The mask condition controller, comprising a mask encoder ($\phi_{\mathrm{mask}}$) and a downsampling module ($\phi_{\mathrm{down}}$), is jointly trained on image-text pairs with and without mask annotations, while all other parameters (including the autoencoder, text encoder, and diffusion UNet) remain frozen.
+The training objective $\mathcal{L}_{c}$ is:
+$\mathcal{L}_{c}=\mathbb{E}_{t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{c}(\mathbf{z}_{t},t,\mathcal{T},\mathcal{M})||_{2}^{2}\Big{]}$
+Here, incorporating data without mask annotations prevents the model from overfitting to those with masks.
 Discussion.
-(i) The large-scale image-text dataset ($\mathcal{D}_{u}$) comprises a wide range of modalities, covering those of source-domain dataset ($\mathcal{D}_{s}$) and target-domain dataset ($\mathcal{D}_{t}$); (ii) Mask-conditioned finetuning on mask-annotated data allows MRGen to learn correlations between images and organ masks; (iii) Such a two-stage training strategy then enables MRGen to extend its controllable generation ability towards target domains lacking mask annotations.
+Such a two-stage training strategy empowers MRGen to achieve controllable generation across diverse modalities, even for those lacking mask annotations, driven by two key factors: (i) text-guided pretraining on large-scale image-text data of various modalities equips MRGen with the foundational knowledge to synthesize diverse MR images based on text prompts; (ii) mask-conditioned finetuning on partial annotated data instructs MRGen to integrate text and mask conditions, enabling controllability that generalizes to modalities included during pretraining.
 Figure 3: Synthetic Data Construction Pipeline.
-MRGen takes text prompt and mask as conditions for controllably generating radiology images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
+MRGen takes text prompt and mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
 3.5 Synthetic Data for Segmentation Training
-With our data engine, we aim to generate target-domain data conditioned on the segmentation masks, to facilitate training segmentation models towards unannotated modalities.
-Data synethesis.
-At inference time, we feed the text prompt ($\mathcal{T^{\prime}}_{t}$) and organ mask ($\mathcal{M^{\prime}}_{t}$) as conditions into our MRGen model ($\Phi_{\mathrm{MRGen}}$) to generate the corresponding target-domain image ($\mathcal{I^{\prime}}_{t}$).
-For simplicity, when constructing the input conditions, we directly utilize region, organ information, and mask annotations from source-domain data.
-Data autofilter.
-To ensure the fidelity of generated images to the mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2 [41] model, as depicted in Figure 3.
-Specifically, we feed the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) into SAM2 to predict a segmentation confidence score ($s_{\mathrm{conf}}$), and a pseudo mask annotation used to calculate the IoU score ($s_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}_{t}$.
-A sample is considered high-quality and aligned with the mask condition if both its IoU score and confidence score exceed the predefined thresholds; otherwise, it is discarded.
-Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [57] CycleGAN [58] MRGen (Ours)
-CM. T1 CM. T2-SPIR 156.98 228.16 157.77 44.82
-CM. T2-SPIR CM. T1 156.98 261.97 188.91 60.79
-MP. T2 MP. ADC 305.56 422.73 112.82 99.38
-MP. ADC MP. T2 305.56 416.31 190.82 123.55
-PS. T1 PS. T2 76.95 120.36 237.52 34.35
-PS. T2 PS. T1 76.95 126.27 89.26 58.90
-LQ. T1 CM. T2-SPIR 109.46 281.56 182.62 88.76
-CM. T2-SPIR LQ. T1 109.46 246.73 260.06 106.45
-MP. ADC PR. T2 387.29 434.54 221.27 116.35
-PR. T2 MP. ADC 387.29 365.10 140.72 88.43
-AM. CT CM. T1 148.04 226.23 248.06 42.53
-AM. CT CM. T2-SPIR 164.73 259.96 268.49 65.30
-ML. CT CM. T2-SPIR 176.51 265.64 307.79 96.72
-Average FID $\downarrow$ 197.06 281.20 200.47 78.95
-Table 2: Quantitative Results (FID) on Generation.
-Smaller values indicate better performance.
-Here, CM., MP., PS., LQ., PR., AM., and ML., denote CHAOS-MRI, MSD-Prostate, PanSeg, LiQA, PROMISE12, AMOS22, and MSD-Liver, respectively.
-Source Dataset Source Modality Target Dataset Target Modality DualNorm [57] nnUNet [21] UMamba [34]
-$\mathcal{D}_{s}$ CycleGAN MRGen $\mathcal{D}_{s}$ CycleGAN MRGen Oracle Box SAM2 nnUNet [21] $\mathcal{D}_{t}$
-Same Datasets, Different Modalities
-CHAOS-MRI T1 CHAOS-MRI T2-SPIR 14.00 6.90 7.58 66.18 4.02 10.58 67.35 45.45 53.12 83.90
-CHAOS-MRI T2-SPIR CHAOS-MRI T1 12.50 0.80 1.38 58.10 0.62 0.22 57.24 43.48 51.94 90.60
-MSD-Prostate T2 MSD-Prostate ADC 1.43 5.52 40.92 57.83 0.19 45.06 52.58 61.50 65.39 82.35
-MSD-Prostate ADC MSD-Prostate T2 12.94 22.20 57.06 61.95 11.80 62.00 64.05 61.07 66.40 89.80
-PanSeg T1 PanSeg T2 0.21 0.68 2.40 9.78 0.38 2.13 9.34 31.69 41.11 82.28
-PanSeg T2 PanSeg T1 0.11 0.30 3.59 12.07 0.27 5.08 10.29 30.94 41.18 80.27
-Different Datasets, Different Modalities
-LiQA T1 CHAOS-MRI T2-SPIR 19.23 16.20 7.84 31.73 11.05 8.65 37.30 62.72 70.86 70.83
-CHAOS-MRI T2-SPIR LiQA T1 31.09 15.80 41.02 57.28 10.33 41.22 52.54 56.11 67.99 93.60
-MSD-Prostate ADC PROMISE12 T2 1.43 17.19 42.13 35.33 23.71 43.24 37.12 50.41 53.61 87.00
-PROMISE12 T2 MSD-Prostate ADC 9.84 19.20 59.95 56.88 21.75 57.21 49.77 61.50 65.39 82.35
-Different Datasets, Different Modalities, CT - MRI
-AMOS22 CT CHAOS-MRI T1 19.78 0.11 6.75 22.50 0.05 8.06 26.73 43.48 51.94 90.60
-AMOS22 CT CHAOS-MRI T2-SPIR 16.09 8.88 52.49 56.23 3.19 43.21 60.53 45.45 53.12 83.90
-MSD-Liver CT CHAOS-MRI T2-SPIR 1.58 3.12 10.14 38.67 1.65 11.06 40.93 62.72 70.86 70.83
-Average DSC score 10.79 8.99 25.63 43.43 6.85 25.89 43.52 50.50 57.92 83.72
+With our data engine, we can then produce MR samples for training downstream segmentation models.
+At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our MRGen model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$).
+To ensure the fidelity of the generated images on mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2-Large [48] model, as depicted in Figure 3.
+Specifically, we feed the conditional mask ($\mathcal{M^{\prime}}$) and the generated image ($\mathcal{I^{\prime}}$) into SAM2 to predict a segmentation map with a confidence score ($s_{\mathrm{conf}}$), which is used to calculate the IoU score ($s_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}$.
+A sample is considered to be high-quality and aligned with the mask condition if both its IoU score ($s_{\mathrm{IoU}}$) and confidence score ($s_{\mathrm{conf}}$) exceed the predefined thresholds; otherwise, it is discarded.
+4 Experiments
+Here, we first outline the experimental settings in Sec.
+4.1, followed by a comprehensive evaluation from both quantitative and qualitative perspectives in Sec.
+4.2 and Sec.
+4.3.
+Lastly, we present ablation study results in Sec.
+4.4 to prove the effectiveness of our proposed method.
+4.1 Experimental Settings
+Unlike existing work that focuses on data augmentation for well-annotated modalities, we explore the more challenging scenarios where certain modalities lack annotations entirely, and aim to leverage generative models to synthesize data for training segmentation models towards these underrepresented modalities.
+Specifically, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within our MRGen-DB, each comprising a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$).
+Models trained on each dataset pair synthesize target-domain samples for training segmentation models.
+We assess our data engine from two aspects: (i) image generation quality and (ii) downstream segmentation performance on the target-domain test set.
+Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [68] CycleGAN [69] UNSB [27] MRGen (Ours)
+CM. T1 CM. T2-SPIR 156.98 228.16 157.77 160.12 44.82
+CM. T2-SPIR CM. T1 156.98 261.97 188.91 141.15 60.79
+MP. T2 MP. ADC 305.56 422.73 112.82 303.19 99.38
+MP. ADC MP. T2 305.56 416.31 190.82 346.60 123.55
+PS. T1 PS. T2 76.95 120.36 237.52 80.76 34.35
+PS. T2 PS. T1 76.95 126.27 89.26 76.91 58.90
+LQ. T1 CM. T2-SPIR 109.46 281.56 182.62 193.05 88.76
+CM. T2-SPIR LQ. T1 109.46 246.73 260.06 192.80 106.45
+MP. ADC PR. T2 387.29 434.54 221.27 200.55 116.35
+PR. T2 MP. ADC 387.29 365.10 140.72 252.64 88.43
+Average FID $\downarrow$ 207.25 290.37 178.18 194.78 82.18
+Table 2: Quantitative Results (FID) on Generation. Here, CM., MP., PS., LQ., and PR., denote CHAOS-MRI, MSD-Prostate, PanSeg, LiQA, and PROMISE12, respectively.
+Baselines.
+For generation, we compare generated images from MRGen against three representative approaches: CycleGAN [69] and UNSB [27] for translating source-domain images to the target domain; and DualNorm [68] for exhaustive augmentation of source-domain images.
+For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data.
+We adopt nnUNet [23] and UMamba [40] as segmentation frameworks for all settings, except for DualNorm, which employs a customized UNet following their official codes.
+Comparisons with additional baselines are provided in Sec.
+D.2.
+Evaluation metrics.
+We employ distinct metrics for the evaluation: For image generation, we employ Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images.
+For segmentation models, we employ commonly used Dice Similarity Coefficient (DSC) [41] score to compare predicted masks with ground truth.
+Considering segmentation consistency, we stack slices into 3D volumes, calculate the DSC for each organ individually, and average them as the final result.
+Implementation details.
+All images are resized to 512 $\times$ 512, and training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [38] optimizer.
+We start by training the autoencoder with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations.
+Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively.
+Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance [19].
+The compression ratio, latent dimension $d$, KL loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively.
+During inference, we perform 50-step sampling using DDIM [53], with classifier-free guidance weight $w=7.0$.
+For each mask, we generate 20 image candidates and select the best two satisfying the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively.
+Conditions for target-domain data synthesis are directly derived from region, organs information, and segmentation masks of source-domain data.
+Source Dataset Source Modality Target Dataset Target Modality DualNorm [68] UMamba [40] nnUNet [23]
+$\mathcal{D}_{s}$ CycleGAN UNSB MRGen $\mathcal{D}_{s}$ CycleGAN UNSB MRGen
+CHAOS-MRI T1 CHAOS-MRI T2-SPIR 14.00 4.02 10.58 20.56 67.35 6.90 7.58 14.03 66.18
+CHAOS-MRI T2-SPIR CHAOS-MRI T1 12.50 0.62 0.22 4.41 57.24 0.80 1.38 6.44 58.10
+MSD-Prostate T2 MSD-Prostate ADC 1.43 0.19 45.06 45.77 52.58 5.52 40.92 52.99 57.83
+MSD-Prostate ADC MSD-Prostate T2 12.94 11.80 62.00 36.47 64.05 22.20 57.06 38.39 61.95
+PanSeg T1 PanSeg T2 0.21 0.38 2.13 3.08 9.34 0.68 2.40 2.38 9.78
+PanSeg T2 PanSeg T1 0.11 0.27 5.08 4.46 10.29 0.30 3.59 6.68 12.07
+LiQA T1 CHAOS-MRI T2-SPIR 19.23 11.05 8.65 5.14 37.30 16.20 7.84 4.79 31.73
+CHAOS-MRI T2-SPIR LiQA T1 31.09 10.33 41.22 28.52 52.54 15.80 41.02 15.28 57.28
+MSD-Prostate ADC PROMISE12 T2 1.43 23.71 43.24 43.64 37.12 17.19 42.13 42.40 35.33
+PROMISE12 T2 MSD-Prostate ADC 9.84 21.75 57.21 49.82 49.77 19.20 59.95 55.13 56.88
+Average DSC score 10.28 8.41 27.54 24.19 43.76 10.48 26.39 23.85 44.71
 Table 3: Quantitative Results (DSC score) on Segmentation.
-Oracle Box denotes the DSC score compared between the perturbed oracle box prompts for SAM2, and the ground truth; $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$ denote training with real and manually annotated source-domain and target-domain data respectively.
-Results are organized by three different cross-modality simulations, where the best results under nnUNet and UMamba are bolded respectively, while the second best are underlined.
-5 Experiments
-Here, we start by describing the experimental settings in Section 5.1; then, in Section 5.2 and 5.3, we present a comprehensive evaluation of our method from both quantitative and qualitative perspectives; lastly, we provide results for ablation experiments in Section 5.4 to prove the effectiveness of our proposed methods.
-5.1 Experimental Settings
-We aim to evaluate our proposed data engine from two aspects: (i) image generation, and (ii) segmentation.
-Concretely, to simulate MRI segmentation towards unannotated modalities, we construct 8 cross-modal dataset pairs within our MedGen-1M.
-Each pair comprises a mask-annotated source-domain dataset and an unannotated target-domain dataset.
-The generative model is trained on each dataset pair to synthesize target-domain samples, which are then used to train segmentation models.
-We assess the quality of generated images, and the performance of the segmentation models on the target-domain test set.
-Baselines.
-For image generation, we compare mask-conditioned generated images from MRGen against two other approaches: CycleGAN [58] for translating source-domain images to the target domain; and DualNorm [57] for aggressive augmentation of source-domain images.
-For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) MRGen generated data; and (v) manually annotated target-domain data, serving as a strong oracle baseline.
-We adopt nnUNet [21] and UMamba [34] as segmentation frameworks for all settings, except for DualNorm, which uses a customized UNet.
-Moreover, we include SAM2 [41] as a reference for interactive semi-automatic segmentation, using randomly perturbed oracle boxes as prompts.
-Evaluation metrics.
-We employ different metrics for assessment: For image generation, we adopt the commonly used Fréchet Inception Distance score (FID) [16] to assess the diversity and quality of the generated images.
-For the downstream segmentation models, we employ the widely used Dice Similarity Coefficient (DSC) [35] to compare the predicted segmentation mask and ground truth.
-Considering segmentation consistency, we stack slices back into 3D volumes for DSC calculation.
+Here, $\mathcal{D}_{s}$ denotes training with manually annotated source-domain data.
+Results with the best and second best results are bolded and underlined, respectively.
 Figure 4: Qualitative Results of Controllable Generation.
-We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference.
-Here, prostate, liver, right kidney, left kidney and spleen are contoured with different colors.
-Implementation details.
-All the images are resized to resolution 512 $\times$ 512 pixels.
-For training, we start by training the VAE with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations.
-Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively.
-Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance for better performance.
-All training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [32] optimizer.
-The downsampling factor, latent feature dimension $d$, KL divergence loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively.
-During inference, we perform 50-step sampling using DDIM [45], with the classifier-free guidance [17] weight set to $w=7.0$.
-We adopt the pretrained text encoder in BiomedCLIP [52] to encode text prompts, and SAM2 [41] for automatic filtering.
-For each conditional mask, we generate 20 candidates and select the best two that satisfy the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively.
-Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities.
-We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth.
-We visualize prostate on the middle row, and liver, right kidney, left kidney and spleen on the upper and lower row with different colors.
+We present images from source domains ($\mathcal{D}_{s}$) and target domains ($\mathcal{D}_{t}$) for reference.
+Here, liver, right kidney, left kidney, spleen, prostate, and pancreas are contoured with different colors.
 4.2 Quantitative Results
-Image generation.
-As shown in Table 2, images from the source domain exhibit high FID values with the target domain, indicating substantial discrepancies between distinct radiology modalities.
-Similarly, images augmented by DualNorm and translated by CycleGAN also exhibit high FID values, confirming their limited ability to emulate target-domain images.
-Conversely, our MRGen presents a significantly lower FID, demonstrating its capability to accurately generate images in target modalities, thus providing a foundation for training segmentation models with synthetic data.
+Generation.
+As shown in Table 2, source-domain images exhibit high FID values compared to the target domain, indicating substantial discrepancies across distinct modalities.
+Similarly, images augmented by DualNorm and translated by CycleGAN and UNSB also show high FID values, confirming their limited ability to emulate target-domain images.
+In contrast, our MRGen presents a significantly lower FID, demonstrating its ability to accurately generate images of target modalities, providing a foundation for training segmentation models with high-quality synthetic data.
 Segmentation.
-As presented in Table 3, we have the following four key observations: (i) The significant discrepancies among different modalities hinder the ability of nnUNet and UMamba trained exclusively on source-domain data to generalize to the target domain, resulting in the lowest average DSC scores.
-(ii) The conventional image translation approach, CycleGAN, and the augmentation-based method, DualNorm, achieve moderate improvements in average DSC scores.
-However, these methods still struggle in the target domain and occasionally underperform in specific scenarios.
-(iii) In contrast, our proposed MRGen generates high-quality images for target domains, achieving the highest DSC score in 11 out of 13 experiments and significantly outperforming CycleGAN.
-Notably, MRGen consistently improves performance across all settings compared to source-only training, demonstrating the adaptability and effectiveness of its synthetic data across various segmentation architectures.
-(iv) The interactive SAM2 model, trained on substantially larger datasets, exhibits robust zero-shot segmentation capabilities with perturbed oracle box prompts.
-However, as a semi-automatic method, its reliance on spatial prompts and manual intervention limits its scalability and practical applicability.
+As depicted in Table 3, we draw the following three observations: (i) significant discrepancies between different modalities challenge the generalization ability of nnUNet and UMamba models trained solely on source-domain data, leading to notably lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements in average DSC scores, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the highest DSC score in 8 out of 10 experiments, significantly outperforming others.
+Notably, MRGen consistently improves performance of nnUNet and UMamba, demonstrating the adaptability and robustness of its synthetic data across various segmentation architectures.
+More comparisons will be included in the Sec.
+D.2.
 4.3 Qualitative Results
 Generation.
-As depicted in Figure 4, the images of distinct modalities exhibit substantial visual discrepancies, making it challenging for DualNorm to simulate via augmentation.
-Additionally, CycleGAN is prone to instability during training, frequently leading to model collapse when learning such complex translations.
-Conversely, our proposed MRGen effectively generates images that closely resemble the target domain and align with conditioned organ masks.
-This capability provides a solid foundation for the controlled generation of training data to extend segmentation models towards unannotated modalities.
+As presented in Figure 4, images of distinct modalities exhibit substantial visual discrepancies, making it challenging for DualNorm to simulate via exhaustive augmentation.
+While CycleGAN and UNSB preserve contours well, they suffer from unstable training and model collapse when learning complex transformations, thus failing to synthesize target-domain images accurately.
+In contrast, MRGen effectively generates images that closely resemble target domains and align with given organ masks, providing compelling evidence for controllable MRI data synthesis.
+Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities.
+Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging.
+We visualize liver, right kidney, left kidney and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
 Segmentation.
-As shown in Figure 5, despite significant variability between the source and target domains, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data.
-However, the data derived from DualNorm and CycleGAN lead to unsatisfactory segmentation results, due to notable discrepancies from the target domain or low quality.
+As illustrated in Figure 5, despite significant appearance variations among distinct modalities, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data.
+However, DualNorm, and segmentation models trained on data derived from CycleGAN and UNSB, yield unsatisfactory results.
+Model SDM [49] SDM-ft MRGen-M MRGen
+PSNR $\uparrow$ 31.32 35.65 42.62
+SSIM $\uparrow$ 0.989 0.996 0.999
+MSE $\downarrow$ 0.0037 0.0014 0.0003
+FID $\downarrow$ 249.24 91.48 41.82 39.63
+CLIP-I $\uparrow$ 0.3151 0.6698 0.7512 0.8457
+CLIP-T $\uparrow$ 0.1748 0.3199 0.3765 0.3777
+Table 4: Ablation on Reconstruction and Text-guided Generation. Here, MRGen-M adopts the same autoencoder as MRGen.
+4.4 Ablation Studies
+To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation task, as follows.
+Generation.
+We assess autoencoder reconstruction and text-guided generation performance on MRGen-DB test set across various models, including: (i) pretrained Stable Diffusion (SDM), (ii) SDM finetuned on MRGen-DB (SDM-ft), (iii) our model conditioned only on free-text modality labels (MRGen-M), and (iv) our full MRGen with text prompts, incorporating modalities, attributes, regions, and organs information.
+The reconstruction quality is assessed using PSNR, SSIM, and MSE between the reconstructed and original images.
+For generation, we use FID, along with image-to-image (CLIP-I) and image-to-text (CLIP-T) similarities between generated images and ground truth or modality labels, computed by BiomedCLIP [62].
+As presented in Table 4, we can draw the following observations: (i) while SDM pretrained on natural images performs poorly on MRI, finetuning on MRGen-DB yields substantial improvement in both reconstruction and generation; (ii) MRGen, with higher latent dimensions (16 versus 4 of SDM) and the BiomedCLIP text encoder, achieves significantly better performance; and (iii) our templated text prompts further enable MRGen to better distinguish distinct modalities, regions, and organs, leading to superior synthesis quality.
+To this end, we employ high-capacity autoencoders trained on MRGen-DB, text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
 Method AutoFilter $\mathcal{D}_{t}$ Image CHAOS-MRI [26] MSD-Prostate [2]
@@
 66.18 58.10 57.83 61.95
-Table 4: Ablation Study on Segmentation Performance (DSC score). Here, T2S. denotes T2-SPIR.
-5.4 Ablation Studies
-To validate the effectiveness of our modules and strategies, we conduct ablation studies on the downstream segmentation task.
-Specifically, we train nnUNet using data synthesized under different training and inference settings.
-As presented in Table 4, we have the following observations: (i) even without integrating unannotated target-domain data during training, MRGen still improves segmentation performance on the target modality.
-This indicates that the two-stage training strategy effectively enables MRGen to extend controllable generation to modalities without mask annotations; (ii) the autofilter pipeline based on SAM2 further improves performance by selecting data pairs that align better with the mask conditions.
-More ablation studies on image generation will be provided in Section D.2 of the Appendix.
+Table 5: Ablation on Segmentation Performance (DSC score).
+Segmentation.
+We train nnUNet with data generated by MRGen under various training and inference settings.
+As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without incorporating target-domain unannotated data during training, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain unannotated images and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
 5 Conclusion
-This paper explores a novel paradigm of applying generative models in medical imaging: controllably synthesizing data for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios.
-To support this, we curate a large-scale radiology image-text dataset, MedGen-1M, featuring detailed modality labels, attributes, regions, and organ information, along with a subset of organ mask annotations.
-Built upon the dataset, our diffusion-based controllable data engine, MRGen, generates MR images conditioned on text prompts and masks, enabling data synthesis across unannotated modalities for training segmentation models.
-Comprehensive quantitative and qualitative evaluations across diverse modalities demonstrate that MRGen can effectively synthesize data, advancing MRI segmentation for unannotated modalities.
+This paper explores generative models for controllable MRI generation, particularly to facilitate training segmentation models for underrepresented modalities lacking mask annotations.
+To support this, we curate a large-scale radiology image-text dataset, MRGen-DB, featuring detailed modality labels, attributes, regions, and organ information, with a subset of organ mask annotations.
+Built on this, our diffusion-based data engine, MRGen, synthesizes MR images of various annotation-scarce modalities conditioned on text prompts and masks.
+Comprehensive evaluations across diverse modalities demonstrate that MRGen effectively improves segmentation performance on unannotated modalities by producing high-quality synthetic data.
+We believe this will offer new insights into addressing the scarcity of medical data and annotations, holding clinical significance.
 Appendix
@@
 The forward diffusion and reverse denoising processes are then performed on the latent codes ($\mathbf{z}$) within this latent space, significantly reducing computational cost and improving efficiency.
-Appendix B Details of MedGen-1M
-In this section, we provide additional details about our collected and curated MedGen-1M dataset.
-In Section B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Section B.2, we present more comprehensive statistics about the dataset.
+Appendix B Details of MRGen-DB & Synthetic Data
+In this section, we provide additional details about our collected and curated MRGen-DB dataset.
+In Sec.
+B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Sec.
+B.2, we present more comprehensive dataset statistics.
+Additionally, in Sec.
+B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training under each experimental setting.
 B.1 Automatic Annotations
-We employ an automated annotation pipeline to annotate the data in our MedGen-1M, ensuring that the samples contain comprehensive clinically relevant annotation information.
-This process primarily includes two components: human body region classification and modality explanation, which will be detailed as follows.
+We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and comprehensive clinically relevant information to distinguish distinct modalities, regions, and organs.
+This process primarily consists of two precise and controllable components: human body region classification and modality explanation, which will be detailed as follows.
 Region classification.
@@
 Now, tell me the attributes of \$modality\$.
+To ensure reliability and accuracy, we have randomly and uniformly sampled approximately 2% (5K out of 250K) of region annotations and 20% (60 out of $\sim$300) of modality attribute annotations for manual verification, achieving high accuracies of 95.33% and 91.67%.
+Furthermore, the effectiveness in downstream tasks also validates the quality of automatic annotations.
 B.2 Dataset Statistics
-In this section, we present more detailed statistics about our curated MedGen-1M dataset, including the unannotated image-text pairs from Radiopaedia333radiopaeida.org, as well as the mask-annotated data sourced from various open-source datasets.
+In this section, we present more detailed statistics about our curated MRGen-DB dataset, including the unannotated image-text pairs from Radiopaedia333radiopaeida.org, as well as the mask-annotated data sourced from various open-source datasets.
+Figure 6: Data Statistics of Radiopaedia-MRI.
+(a) Distribution of slice counts across various modalities in Radiopaedia-MRI; (b) Proportional distribution of slices across different regions in Radiopaedia-MRI.
 Data without mask annotations.
-For the image-text pairs from Radiopaedia-CT and Radiopaedia-MRI, which are used for training the autoencoder and text-guided generation, we allocate 1% of the data as a test set to evaluate reconstruction and generation performance, maximizing the amount of data available for pretraining.
-As a result, 1,007,616 samples are used for training, comprising 804,628 from Radiopaedia-CT and 202,988 from Radiopaedia-MRI.
-The test set consists of 10,179 samples, with 8,128 from Radiopaedia-CT and 2,051 from Radiopaedia-MRI.
+For the image-text pairs from Radiopaedia-MRI, which are used for training the autoencoder and text-guided generation, we allocate 1% of the data as a test set to evaluate reconstruction and generation performance, maximizing the amount of data available for pretraining.
+As a result, 202,988 samples are used for training, and the test set consists of 2,051 samples.
 We conduct a statistical analysis of the distribution of modalities in Radiopaedia-MRI, as presented in Figure 6 (a).
 The free-text modality labels cover approximately 300 categories, providing a diverse set of MRI modalities that form a crucial foundation for MRGen to learn text-guided generation and expand its mask-conditioned generation capabilities towards modalities originally lacking mask annotations.
-Furthermore, the distribution of images across different regions in Radiopaedia-CT and Radiopaedia-MRI is presented in Figure 6 (b).
-Figure 6: Data Statistics of Radiopaedia.
-(a) Distribution of slice counts across various modalities in Radiopaedia-MRI; (b) Proportional distribution of slices across different regions in Radiopaedia-CT and Radiopaedia-MRI.
+Furthermore, the distribution of images across different regions in Radiopaedia-MRI is presented in Figure 6 (b).
 Data with mask annotations.
@@
 # Vol. # Slc. # Slc. w/ mask # Vol. # Slc. # Slc. w/ mask
-LiQA [31] Liver T1-weighted 24 1,718 1,148 6 467 298
 PanSeg [65] Pancreas T1-weighted 309 14,656 5,961 75 3,428 1,400
@@
 ADC 26 492 100 6 110 83
-PROMISE12 [29] Prostate T2-weighed 40 1,137 645 10 110 133
 CHAOS-MRI [26] Liver, Right Kidney, T1-weighted 32 1,018 770 8 276 230
 Left Kidney, Spleen T2-SPIR 16 503 388 4 120 104
-MSD-Liver [2] Liver CT 105 46,695 15,260 26 11943 3903
-AMOS22CT [23] Liver, Right Kidney, CT 200 26,069 19,023 100 15,361 11,256
-Left Kidney, Spleen, etc.*
-Total / / 1,083 105,074 48,501 218 34,907 18,802
-Table 5: Details of Segmentation-annotated Datasets in MedGen-1M.
+PROMISE12 [34] Prostate T2-weighed 40 1,137 645 10 240 133
+LiQA [36] Liver T1-weighted 24 1,718 1,148 6 467 298
+Total / / 778 36,710 14,218 192 7,733 3,643
+Table 6: Details of Segmentation-annotated Datasets in MRGen-DB.
 Here, # Vol.
@@
 w/ mask indicates the number of 2D slices with mask annotations.
-*: AMOS22CT [23] also contains images and mask annotations for the following organs: Gallbladder, Esophagus, Stomach, Aorta, Inferior Vena Cava, Pancreas, Duodenum, Urinary Bladder, and Adrenal Gland.
+B.3 Synthetic Data Statistics
+This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings.
+Concretely, we use mask annotations from the entire source-domain dataset (including both training and test sets) as input conditions to generate target-domain images, forming image-mask training pairs.
+Exceptions include: (i) for the MSD-Prostate [2] dataset, where images of T2 and ADC modalities have already been registered, we restrict inputs to the source-domain training set to prevent data leakage; and (ii) for dataset pairs with CHAOS-MRI-T1 [26] as the target domain, each source-domain mask is used twice to synthesize both T1-InPhase and T1-OutofPhase data.
+This setup is consistent across all baselines.
+Additionally, with our proposed autofilter pipeline, MRGen generates 20 candidate images per mask and selects the top two that meet the predefined thresholds.
+If no samples satisfy the thresholds, all thresholds will be relaxed by 0.10, and the sample of the highest quality is chosen, ensuring full exploitation of source-domain masks.
+Otherwise, all low-quality generated samples are discarded to avoid noisy data.
+Finally, the statistics of target-domain training pairs generated by MRGen for each experimental setting are summarized in Table 7.
+Source Dataset Source Modality Target Dataset Target Modality # Slices ($\mathcal{D}_{s}$) # Synthetic Data
+CHAOS-MRI T1 CHAOS-MRI T2-SPIR 1,294 433
+CHAOS-MRI T2-SPIR CHAOS-MRI T1 607 1,118
+MSD-Prostate T2 MSD-Prostate ADC 492 775
+MSD-Prostate ADC MSD-Prostate T2 492 745
+PanSeg T1 PanSeg T2 18,084 2,160
+PanSeg T2 PanSeg T1 15,276 2,215
+LiQA T1 CHAOS-MRI T2-SPIR 2,185 2,267
+CHAOS-MRI T2-SPIR LiQA T1 607 636
+MSD-Prostate ADC PROMISE12 T2 602 742
+PROMISE12 T2 MSD-Prostate ADC 1,377 1,077
+Table 7: Synthetic Data Statistics.
+Here, # Slices ($\mathcal{D}_{s}$) denotes the number of source-domain samples under each experimental setting, which serves as input for translation-based baselines.
+Moreover, # Synthetic Data represents the total volume of data generated by MRGen.
 Appendix C Implementation Details
 In this section, we will provide a comprehensive explanation of the implementation details discussed in the paper.
-Specifically, Section C.1 describes the preprocessing and augmentation strategies applied to the training data.
-Section C.2 elaborates on the details of the autofilter pipeline.
-Finally, Section C.3 outlines the implementation details of the baselines.
+Concretely, Sec.
+C.1 describes the preprocessing and augmentation strategies applied to the training data.
+Sec.
+C.2 elaborates on the details of the autofilter pipeline.
+Finally, Sec.
+C.3 outlines the implementation details of various baselines.
 C.1 Preprocessing & Augmentation
 Data preprocessing.
-To ensure consistency across data from various sources and modalities, we apply tailored preprocessing strategies as follows: (i) For data from Radiopaedia, the images are directly rescaled to the range [0, 1]; (ii) For CT images with mask annotations, intensities are clipped to [-300, 200] and then rescaled to [0, 1]; (iii) For MR images with mask annotations, intensities are clipped to the 0.5 and 99.5 percentiles and rescaled to [0, 1].
+To ensure consistency across data from various sources and modalities, we apply tailored preprocessing strategies as follows: (i) For data from Radiopaedia-MRI, the images are directly rescaled to the range [0, 1]; (ii) For MR images with mask annotations, intensities are clipped to the 0.5 and 99.5 percentiles and rescaled to [0, 1].
 After normalization, all data are subsequently rescaled to [-1, 1] for training various components of MRGen, including the autoencoder, diffusion UNet, and mask condition controller.
-For training downstream segmentation models, images are rescaled to [0, 255] and saved in ‘.png’ format, followed by the preprocessing configurations of nnUNet [21] and UMamba [34].
+For training downstream segmentation models, images are rescaled to [0, 255] and saved in ‘.png’ format, followed by the official preprocessing configurations of nnUNet [23] and UMamba [40].
 Data augmentation.
-During autoencoder training, we apply random data augmentations to the images with a 20% probability.
-These augmentations included horizontal flipping, vertical flipping, and rotations of 90, 180, 270.
+During autoencoder training, we apply random data augmentations to images with a 20% probability.
+These augmentations include horizontal flipping, vertical flipping, and rotations of 90, 180, 270.
 In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller.
+We do not explicitly adopt data balancing, as we empirically find that it does not lead to significant performance changes.
 For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [40].
 C.2 Autofilter Pipeline
-When deploying our proposed data engine, MRGen, to synthesize data for segmentation model training, we adopt the off-the-shelf SAM2 [41] model to perform automatic interactive segmentation on generated images, with the mask conditions as prompt.
-This pipeline automatically selects samples that are fidelity to the condition masks, thus ensuring the quality of synthetic image-mask pairs.
-In this section, we elaborate on more implementation details of this automatic filtering pipeline, particularly focusing on the generation of MR images that encompass multiple organs.
+When deploying our proposed data engine, MRGen, to synthesize training data for segmentation models, we adopt the off-the-shelf SAM2-Large [48] model to perform automatic interactive segmentation on generated images, with the mask conditions as spatial prompts.
+Empirically, we observe that SAM2 consistently segments images based on their contours, guided by the provided spatial prompts.
+Concretely, it produces high-quality pseudo mask annotations for images with contours closely matching mask conditions, while performing poorly for synthesized images that deviate significantly from mask conditions.
+This characteristic allows our pipeline to automatically filter out samples faithful to the condition masks and discard erroneous ones, thus ensuring the quality of synthesized image-mask pairs.
+Here, we elaborate on more implementation details of this automatic filtering pipeline, particularly focusing on the generation of MR images that encompass multiple organs.
 Specifically, we begin by defining the following thresholds: confidence threshold ($\tau_{\mathrm{conf}}$), IoU score threshold ($\tau_{\mathrm{IoU}}$), average confidence threshold ($\bar{\tau}_{\mathrm{conf}}$), and average IoU threshold ($\bar{\tau}_{\mathrm{IoU}}$).
-Both the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) are fed into SAM2.
-For each organ mask $\mathcal{M^{\prime}}^{i}_{t}$ in $\mathcal{M^{\prime}}_{t}$, SAM2 will output a confidence score ($s^{i}_{\mathrm{conf}}$) and a pseudo mask annotation, which is then used to calculate the IoU score ($s^{i}_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}^{i}_{t}$.
-For each generated sample ($\mathcal{I^{\prime}}_{t}$), it is regarded as a high-quality sample if the following conditions are satisfied: $\{s^{i}_{\mathrm{IoU}}\geq\tau_{\mathrm{IoU}},s^{i}_{\mathrm{conf}}\geq\tau_{\mathrm{conf}}\,|\,\forall i\}$, and $\{\bar{s}_{\mathrm{IoU}}\geq\bar{\tau}_{\mathrm{IoU}}$, $\bar{s}_{\mathrm{conf}}\geq\bar{\tau}_{\mathrm{conf}}\}$.
+Both the controlling mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) are fed into SAM2.
+For each organ mask $\mathcal{M^{\prime}}^{i}_{t}$ in $\mathcal{M^{\prime}}_{t}$, SAM2 will output a segmentation map with a confidence score ($s^{i}_{\mathrm{conf}}$), which is then used to calculate the IoU score ($s^{i}_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}^{i}_{t}$.
+For each generated sample ($\mathcal{I^{\prime}}_{t}$), it is regarded to be high-quality and aligned with the mask condition if the following conditions are satisfied: $\{s^{i}_{\mathrm{IoU}}\geq\tau_{\mathrm{IoU}},s^{i}_{\mathrm{conf}}\geq\tau_{\mathrm{conf}}\,|\,\forall i\}$, and $\{\bar{s}_{\mathrm{IoU}}\geq\bar{\tau}_{\mathrm{IoU}}$, $\bar{s}_{\mathrm{conf}}\geq\bar{\tau}_{\mathrm{conf}}\}$.
 Otherwise, the sample will be discarded.
-In all experiments, the predefined thresholds are set as follows: $\tau_{\mathrm{IoU}}=0.70$, $\tau_{\mathrm{conf}}=0.80$, $\bar{\tau}_{\mathrm{IoU}}=0.80$, and $\bar{\tau}_{\mathrm{conf}}=0.90$.
+For each conditional mask, we synthesize 20 image candidates and select the best two that satisfy the predefined thresholds.
+Across all experiments, the thresholds are set as follows: $\tau_{\mathrm{IoU}}=0.70$, $\tau_{\mathrm{conf}}=0.80$, $\bar{\tau}_{\mathrm{IoU}}=0.80$, and $\bar{\tau}_{\mathrm{conf}}=0.90$.
 C.3 Baselines
-CycleGAN.
-We follow the official implementation and training strategies across all experimental settings.
+In this section, we introduce the implementation details of representative baselines and discuss other relevant methods by category.
+Concretely, we first consider the most related ones, including augmentation-based and translation-based methods.
+Augmentation-based methods.
+These approaches [68, 22, 54, 44, 59, 7] typically rely on mixing multi-domain training data or employing meticulously designed data augmentation strategies.
+Here, we consider the representative one, DualNorm [68].
+Following its official implementation, we apply random non-linear augmentation on each source-domain image, to generate a source-dissimilar training sample, and train the dual-normalization model.
+All preprocessing steps, network architectures, and training strategies adhere to the official recommendations, with the exception that images are resized to 512 $\times$ 512, consistent with other methods.
+Notably, we evaluate DualNorm on all slices in the test set, offering a more rigorous evaluation compared to the official code, which only considers slices with segmentation annotations.
+Translation-based methods.
+These methods [52, 27, 46, 28] are commonly inspired by CycleGAN [69]; therefore, we compare with open-source CycleGAN [69], UNSB [27], and MaskGAN [46].
+We follow their official implementations and training strategies across all experimental settings.
 Subsequently, source-domain images are translated into the target domain and paired with the source-domain masks to create paired samples for training downstream segmentation models.
-SAM2.
-We use SAM2-Large in all the experiments.
+Moreover, we have also explored other approaches leveraging the progress of generative models.
+Generation-based methods.
+Existing medical generation models [60, 14, 15, 9, 55] still struggle with complex and challenging abdominal MRI generation.
+For instance, MAISI [14] and Med-DDPM [9] are tailored for CT and brain MRI synthesis, respectively.
+To adapt to our task, we finetune MAISI [14] on our data, as a generation-based baseline.
+Additionally, we consider other methods aimed at addressing our focused challenge, i.e., segmenting MR images of underrepresented modalities lacking mask annotations.
+These include few-shot learning approaches, general-purpose segmentation models, and methods incorporating oracle inputs as performance references.
+Notably, these approaches, to varying degrees, rely on manually annotated target-domain segmentation masks or external datasets.
+Thus, they should be regarded as references only rather than fair comparisons with the aforementioned methods and our MRGen.
+Few-shot methods.
+Specifically, we compare with a few-shot nnUNet [23] (pre-trained on source-domain data and finetuned on 5% target-domain manually annotated data), as well as UniVerSeg [5] with its official implementation and checkpoint.
+General segmentation models.
+We adopt the official code and checkpoint of TotalSegmentor-MRI [8], which has been trained on extensive manually annotated data and diverse modalities, as a strong general-purpose segmentation baseline.
+Models with oracle inputs.
+We include SAM2-Large [48] as a reference for interactive semi-automatic segmentation, using randomly perturbed oracle boxes as prompts.
+To simulate the error introduced by manual intervention, the oracle boxes are randomly shifted at each corner, by up to 8% of the image resolution, following MedSAM [39].
 Segmentation results are derived in a slice-by-slice and organ-by-organ manner: For each slice with mask annotations, we simulate box prompts for each annotated organ individually.
-To take into account the error introduced by manual intervention, the oracle boxes are randomly shifted at each corner, by up to 8% of the image resolution, following the protocol in MedSAM [33].
-DualNorm.
-Following the official implementation, we apply random non-linear augmentation on each image in the source domain, to generate a source-dissimilar training sample, and train the dual-normalization model.
-All preprocessing steps, network architectures, and training strategies follow the official recommendations, with the exception that images are rescaled to 512 $\times$ 512, consistent with other approaches.
-Note that we evaluate DualNorm on all the slices in the test set, offering a more rigorous evaluation compared to the official implementation, which only considers slices with segmentation annotation.
+Finally, we also include nnUNet [23] trained exclusively on the target-domain mask-annotated dataset ($\mathcal{D}_{t}$) as an oracle reference, reflecting the performance upper bound with sufficient annotated data.
 Appendix D More Experiments
 In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine.
-In Section D.1, we showcase both quantitative and qualitative results of in-domain generation.
-In Section D.2, we conduct extensive ablation studies to illustrate the improvements introduced by our design in radiology image generation.
-Finally, in Section D.3, we provide additional qualitative results to validate the accuracy and flexibility of the generated outputs.
+First, in Sec.
+D.1, we showcase quantitative and qualitative results of in-domain generation.
+Next, in Sec.
+D.2, we present quantitative comparisons with more baselines, further confirming the effectiveness and necessity of our proposed data engine.
+Finally, in Sec.
+D.3, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
 D.1 In-domain Generation
-Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domain.
-Moreover, as presented in Table 6, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated source-domain data.
+Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domains.
+Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated data.
 This offers a feasible solution to address concerns about medical data privacy.
@@
 T2-SPIR T1 83.90 82.06 0.62 57.24 90.60
+Table 8: In-domain & Cross-domain Augmentation Results (DSC score) on Segmentation.
+We compare the performance of nnUNet [23] trained on real data versus synthetic data generated by MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
+Figure 7: Qualitative Results of In-domain Generation.
+Moreover, we provide visualizations of in-domain generation in Figure 7, qualitatively demonstrating that our MRGen can reliably perform controllable generation of a large number of samples within the training domain with mask annotations.
+Dataset Source Modality Target Modality DualNorm nnUNet
+$\mathcal{D}_{s}$ MRGen CycleGAN UNSB MaskGAN MAISI Few-shot UniVerSeg TS-MRI Oracle Box SAM2 nnUNet $\mathcal{D}_{t}$
+CHAOS-MRI T1 T2-SPIR 14.00 6.90 66.18 7.58 14.03 32.73 3.34 52.00 48.91 80.64 45.45 53.12 83.90
+T2-SPIR T1 12.50 0.80 58.10 1.38 6.44 1.89 3.11 53.82 52.79 77.09 43.48 51.94 90.60
+MSD-Prostate T2 ADC 1.43 5.52 57.83 40.92 52.99 29.14 9.15 20.28 0.0 0.0 61.50 65.39 82.35
+ADC T2 12.94 22.20 61.95 57.06 38.39 5.98 6.94 29.38 53.90 0.0 61.07 66.40 89.80
+Average DSC score 10.22 8.86 61.02 26.74 27.96 17.44 5.64 38.87 38.90 39.43 52.88 59.21 86.66
 Table 9: More Quantitative Results (DSC score) on Segmentation.
-We compare the performance of nnUNet [21] trained on real data versus synthetic data generated by our MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
-Moreover, we provide visualizations of in-domain generation in Figure 7, qualitatively demonstrating that our MRGen can reliably perform controllable generation of a large number of samples within the training domain with mask annotations.
-Figure 7: Qualitative Results of In-domain Generation.
-D.2 Ablation Studies
-In this section, we provide more ablation study results on image generation, to demonstrate the necessity of model and data design within our proposed data engine, focusing on autoencoder reconstruction and text-guided generation.
-Autoencoder reconstruction.
-We evaluate the reconstruction performance of various VAEs on the test set derived from our curated MedGen-1M dataset, comprising Radiopaedia-CT and Radiopaedia-MRI.
-The evaluated models include: (i) the VAE of pretrained Stable Diffusion (VAE-SDM), (ii) the VAE of Stable Diffusion finetuned on our dataset (Finetuned-VAE-SDM), (iii) our model trained solely on Radiopaedia-MRI data (VAE-MRGen (MRI-only)), and (iv) our VAE-MRGen trained on the complete dataset (VAE-MRGen (Ours)).
-The reconstruction quality is assessed using standard metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), L1 Error, and Mean Squared Error (MSE) between the reconstructed and original images.
-As illustrated in Table 8, we draw the following three key observations: (i) Finetuning on radiology data: VAEs finetuned on radiology data significantly improve the reconstruction quality compared to pretrained Stable Diffusion models, revealing the limited generalizability of natural image VAEs to radiology images.
-(ii) Latent code dimension: our autoencoders with a larger latent code feature dimension achieve further improvements in reconstruction quality, suggesting that increasing the increased capacity of latent code enables better representations.
-(iii) Extra data: incorporating CT data into training further improves reconstruction quality, approaching the performance of lossless compression.
-To this end, unlike RoentGen [3] that relies on pretrained VAEs, we employ high-capacity autoencoders trained specifically on radiology data, to ensure that the generative models built upon this foundation can produce better image details.
-Text-guided generation.
-We evaluate the text-guided image generation performance of various models on the test set derived from our MedGen-1M dataset, encompassing Radiopaedia-CT and Radiopaedia-MRI.
-The evaluated models include: (i) the pretrained Stable Diffusion (SDM), (ii) the Stable Diffusion finetuned on our MedGen-1M (Finetuned-SDM), (iii) our model using only free-text modality labels as text conditions (MRGen (Modality-only)), and (iv) our MRGen, which utilizes templated text prompts incorporating modality labels, attributes, regions, and organs information (MRGen (Ours)).
-We employ commonly used evaluation metrics for image generation, including the Fréchet Inception Distance (FID [16]) to assess the quality and diversity of generated images.
-Additionally, we utilize BiomedCLIP [52] to calculate two similarity metrics: image-to-image similarity (CLIP-I) between the generated images and ground truth, and image-to-text similarity (CLIP-T) between the generated images and the free-text modality descriptions.
-As presented in Table 8, we draw the following three observations: (i) Finetuning on radiology data: pretrained Stable Diffusion tends to generate natural images rather than ideal radiology images, resulting in suboptimal performance across all evaluation metrics.
-However, finetuning on our dataset significantly enhances its generative performance.
-(ii) Text encoder: using BiomedCLIP [52] as the text encoder, trained on biomedical data, instead of CLIP [40] in Stable Diffusion, improves MRGen’s ability to capture similarities and differences among radiology modalities.
-Additionally, our high-capacity VAE, provides further potential improvements to generation quality.
-(iii) Templated text prompt: incorporating templated text prompts with detailed information (including modality, attributes, regions, and organs) outperforms using modality labels alone.
-This enables MRGen to better grasp the relationships among distinct modalities, thereby further improving generation quality, particularly in terms of image-to-image similarity with the ground truth.
-Model Dim PSNR $\uparrow$ SSIM $\uparrow$ L1 Error $\downarrow$ MSE $\downarrow$
-VAE-SDM [42] 4 31.32 0.989 0.0294 0.0037
-Finetuned-VAE-SDM 4 35.65 0.996 0.0190 0.0014
-VAE-MRGen (MRI-only) 16 40.56 0.998 0.0115 0.0005
-VAE-MRGen (Ours) 16 42.62 0.999 0.0091 0.0003
-Table 7: Ablation Studies on Autoencoder Reconstruction.
-Model FID $\downarrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
-SDM [42] 249.24 0.3151 0.1748
-Finetuned-SDM 91.48 0.6698 0.3199
-MRGen (Modality-only) 41.82 0.7512 0.3765
-MRGen (Ours) 39.63 0.8457 0.3777
-Table 8: Ablation Studies on Text-guided Generation.
+The best and second-best performances are bolded and underlined, respectively.
+Notably, the results marked with a gray background indicate that the corresponding methods may have accessed target-modality annotated data during extensive training (e.g., UniVerSeg, TotalSegmentor-MRI (TS-MRI)), utilized oracle inputs as prompts (e.g., Oracle Box, SAM2), or even been directly trained on target-modality annotated data (e.g., nnUNet (Few-shot), nnUNet ($\mathcal{D}_{t}$)).
+Consequently, these approaches do not represent a fully fair comparison with others, and are primarily included as performance references.
+D.2 More Quantitative Results
+In this section, we compare MRGen with additional baseline methods on two typical cross-modal dataset pairs from MRGen-DB by evaluating the performance of downstream segmentation models, as detailed in the main text.
+Concretely, for both translation-based and generation-based methods, we assess the performance of nnUNet [23] trained on data generated by these methods.
+As depicted in Table 9, we further analyze the relevant baselines by category, as follows.
+Augmentation-based methods.
+Limited to relying on carefully crafted augmentation strategies, DualNorm [68] fails to model nonlinear visual discrepancies among distinct modalities, leading to poor cross-modality segmentation performance.
+Translation-based methods.
+While CycleGAN [69], UNSB [27], and MaskGAN [46] excel at contour preservation, they often suffer from model collapse when learning complex modality transformations, resulting in suboptimal performance.
+Generation-based models.
+Despite finetuned on our dataset, the performance of MAISI [14] is still poor, which we attribute to its lack of modality-conditioning capability.
+This limitation hinders its ability to support cross-modality generation, and consequently, makes it struggle to synthesize target-domain samples for training segmentation models.
+Few-shot methods.
+While few-shot nnUNet [23] and UniverSeg [5] benefit from partial target-domain annotations, MRGen-boosted models outperform without requiring any such annotations, showcasing practical feasibility in clinical scenarios.
+General segmentation models.
+TotalSegmentor-MRI [8] works well on certain datasets/modalities (likely already included during training), but it still performs poorly or even fails on others.
+This significantly limits its practicality in complex clinical scenarios, especially when dealing with underrepresented modalities with diverse imaging characteristics.
+Models with oracle inputs.
+Although SAM2 [48] with perturbed oracle boxes as prompts exhibits impressive zero-shot segmentation capabilities, our MRGen-boosted models still outperform it, trailing only the oracle nnUNet trained directly on target-domain annotated data.
+Moreover, as a semi-automatic method, SAM2’s reliance on high-quality spatial prompts and manual intervention limits its scalability and applicability, while MRGen offers a fully automated, end-to-end solution.
+Overall, MRGen provides a robust, fully automated approach for challenging cross-modality segmentation by producing high-quality synthetic data, with no need for any target-domain mask annotations and proving highly suitable for clinical applications.
+For computational efficiency, we primarily focus on comparing MRGen with some representative baselines, DualNorm [68], CycleGAN [69] and UNSB [27], across more dataset pairs in the main text for a comprehensive evaluation.
 D.3 More Qualitative Results
@@
 Image generation.
-As depicted in Figure 9 and 10, we present additional visualizations of controllable generation on target modalities lacking mask annotations.
+As depicted in Figure 9, we present additional visualizations of controllable generation on target modalities lacking mask annotations.
 These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
 Image segmentation.
-As presented in Figure 11 and 12, we provide more visualizations of segmentation models trained using synthetic data on modalities that originally lack mask annotations.
+As presented in Figure 10, we provide more visualizations of segmentation models trained using synthetic data on modalities that originally lack mask annotations.
 This validates that the samples generated by MRGen can effectively assist in training segmentation models, achieving impressive performance in previously unannotated scenarios.
@@
 E.2 Future Works
-To address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated CT data or collecting more MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
+Due to limited computational resources, we validate our proposed data engine on 2D slices in this paper, with trained segmentation models able to process 3D volumes slice-by-slice.
+However, we believe our idea can be seamlessly extended to 3D volume generation.
+With more computing in the future, we aim to develop a 3D version model for the community, further advancing cross-modality segmentation performance.
+Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
 Figure 9: More Qualitative Results of Controllable Generation.
-We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference.
-Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
-Figure 10: More Qualitative Results of Controllable Generation.
-We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference.
-Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
+We present images from source domains $\mathcal{D}_{s}$ and target domains $\mathcal{D}_{t}$ for reference.
+Here, specific organs are contoured with colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver in LiQA and CHAOS-MRI datasets.
 Figure 10: More Qualitative Results on Segmentation towards Unannotated Modalities.
-We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth.
-Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
-Figure 12: More Qualitative Results on Segmentation towards Unannotated Modalities.
-We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth.
-Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
+Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging.
+Here, specific organs are highlighted with colors: prostate in MSD-Prostate and PROMISE12, pancreas in PanSeg, and liver in LiQA and CHAOS-MRI datasets.

### v3

@@
 Abstract
-Training medical image segmentation models for rare yet clinically significant imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire.
-This paper investigates leveraging generative models to synthesize training data, to train segmentation models for underrepresented modalities, particularly on annotation-scarce MRI.
-Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset having pixelwise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks.
+Training medical image segmentation models for rare yet clinically important imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire.
+This paper investigates leveraging generative models to synthesize data, for training segmentation models for underrepresented modalities, particularly on annotation-scarce MRI.
+Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset featuring pixel-wise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks.
 MRGen can generate realistic images for diverse MRI modalities lacking mask annotations, facilitating segmentation training in low-source domains; (iii) extensive experiments across multiple modalities demonstrate that MRGen significantly improves segmentation performance on unannotated modalities by providing high-quality synthetic data.
@@
 Thus, generative models must enable controllable synthesis based on conditions such as texts, masks, or both.
-Yet, prior works [55, 9, 42, 60, 32] cannot simultaneously support both conditions, limiting their abilitiy to control the generated modalities, regions, and organs effectively.
-Our first contribution is the collection of a large-scale, high-quality radiology image-text dataset, MRGen-DB (short for “Database for MRI Generation”), which includes MRI scans across various modalities sourced from the Internet and open-source repositories.
+Yet, prior works [56, 9, 43, 61, 33] cannot simultaneously support both conditions, limiting their ability to control the generated modalities, regions, and organs effectively.
+Our first contribution is to curate a high-quality, large-scale radiology image-text dataset, termed as MRGen-DB, which includes MRI scans across various modalities sourced from the Internet and open-source repositories.
 The dataset consists of nearly 250,000 2D slices, enriched with detailed annotations such as modality labels, attributes, region, and organ information, with a subset providing organ masks.
-This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training a general MRI generation model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
+This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training an MRI generative model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
 For controllable data generation, we present MRGen, a diffusion-based data engine for MRI synthesis, that supports conditioning on both text prompts and segmentation masks.
-We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) mask-conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks.
-Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enables training segmentation models for these underrepresented modalities with synthetic data.
+We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks.
+Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enabling training segmentation models for the underrepresented modalities with synthetic data.
 Overall, our contributions can be summarized as follows: (i) we explore the use of generative models for MRI synthesis across annotation-scarce modalities, facilitating training segmentation models for underrepresented modalities; (ii) we curate MRGen-DB, a large-scale radiology image-text dataset, which features detailed modality labels, attributes, regions, and organs information, with a subset of organ mask annotations, providing a robust foundation for medical generative modeling; (iii) we develop MRGen, a diffusion-based data engine capable of controllable generation, conditioned on templated text prompts and segmentation masks; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably generate high-quality MR images, improving ‘zero-shot’ segmentation performance for unannotated modalities.
-To our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first general medical generative model tailored for annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
-2 Related works
+To the best of our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first generative model tailored for general annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
+2 Related Work
 Generative models have been a research focus in computer vision for years, with GANs [13] and diffusion models [20, 54] leading the advancements.
@@
 Figure 2: Architecture Overview.
-Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
+Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from MRGen-DB dataset; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
 3 Method
-Here, we start to formulate the problem of interest in Sec.
+We start to formulate the problem of interest in Sec.
 3.1, followed by a detailed description of the dataset curation procedure in Sec.
@@
 where $\Theta$ and $\Theta_{c}$ refer to parameters of the generative model and the mask condition controller, respectively.
-Developing such a controllable data engine, thus enables synthesizing high-quality data to train segmentation models for the challenging ‘zero-shot’ scenario.
+Developing such a controllable data engine enables synthesizing high-quality data to train segmentation models for the challenging scenario that lacks manual annotations, for example, ‘zero-shot’ evaluation.
 Relations to existing tasks.
@@
 Instead, we target the more challenging scenario of synthesizing MR images for scarce and underrepresented modalities where no manual mask annotations are available.
-While image translation methods [69, 27, 52, 46] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data for training or are limited to specific modality conversions.
+While image translation methods [70, 28, 53, 47] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data pairs for training or are limited to specific modality conversions.
 In contrast, our proposed MRGen framework offers a flexible and controllable generation pipeline, enabling the synthesis of complex abdominal MRI data across diverse modalities, even in the absence of mask annotations.
@@
 Automatic annotations.
-Abdominal imaging is highly variable, exhibiting significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region.
+Abdominal imaging normally exhibits significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region.
 Relying solely on modality labels is insufficient to differentiate these distinctions.
 To address this, we divide the abdomen into six anatomical regions: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region.
-Using the pre-trained BiomedCLIP model [62], we automatically categorize all 2D slices into these regions.
-To maintain annotation quality, slices with low confidence scores ($<40\%$) are intentionally left unlabeled.
+Using the pre-trained BiomedCLIP [63], we automatically classify all 2D slices into these regions.
+To maintain annotation quality, slices with low confidence ($<40\%$) are intentionally left unlabeled.
 This process enriches the dataset with detailed region-specific information.
@@
 These detailed descriptions enable the model to understand and differentiate imaging characteristics across modalities.
-To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a subset of the data.
+To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a data subset.
 Specifically, 2% of the region annotations and 20% of the modality attributes have been reviewed, achieving high accuracies of 95.33% and 91.67%, respectively.
@@
 Discussion.
-After completing the data processing above, we assemble the MRGen-DB dataset, which includes approximately 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as presented in Table 1.
-Each sample is paired with its modality label, attributes, region, and organ information, with about 18,000 samples featuring mask annotations.
-The scale, diversity, and fine-grained annotations of MRGen-DB provide sufficient information for training generative models tailored for MR images.
+After the data processing above, we assemble the MRGen-DB dataset, which includes about 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as shown in Table 1.
+In the following parts, we treat each 2D slice as a data sample, which is paired with its modality label, attributes, region, and organ information, with roughly 18,000 samples featuring mask annotations.
+The scale, diversity, and fine-grained annotations of MRGen-DB offer sufficient information for training generative models tailored for MR images.
 More statistics are provided in Sec.
-B.2.
+B.2 of the Appendix.
 3.3 Architecture
@@
 ^$\hat{\mathcal{I}}=\phi_{\mathrm{dec}}(\mathbf{z})=\phi_{\mathrm{dec}}(\phi_{\mathrm{enc}}(\mathcal{I}))$
-To enable the generative model to effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
+To effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
 Text-guided generation.
@@
 As depicted in Figure 2 (b), the denoising process adopts a UNet [51] and reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$.
-Concretely, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”.
+Concretely, to generate images guided by text prompts, we design a templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”.
 These templated prompts ensure sufficient clinical information to distinguish distinct modalities, regions, and organs.
@@
 Autoencoder reconstruction.
-The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [29] and $\gamma$ denotes a predefined weight.
+The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [30], and $\gamma$ denotes a predefined weight.
 Text-guided pretraining.
@@
 Figure 3: Synthetic Data Construction Pipeline.
-MRGen takes text prompt and mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
+MRGen takes a text prompt and a mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
 3.5 Synthetic Data for Segmentation Training
 With our data engine, we can then produce MR samples for training downstream segmentation models.
-At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our MRGen model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$).
+At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our generative model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$).
 To ensure the fidelity of the generated images on mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2-Large [49] model, as depicted in Figure 3.
@@
 4.1 Experimental Settings
-Unlike existing work that focuses on data augmentation for well-annotated modalities, we explore the more challenging scenarios where certain modalities lack annotations entirely, and aim to leverage generative models to synthesize data for training segmentation models towards these underrepresented modalities.
-Specifically, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within our MRGen-DB, each comprising a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$).
-Models trained on each dataset pair synthesize target-domain samples for training segmentation models.
-We assess our data engine from two aspects: (i) image generation quality and (ii) downstream segmentation performance on the target-domain test set.
+Unlike existing work focusing on data augmentation for well-annotated modalities, we explore the more challenging scenario where certain modalities lack annotations entirely, and aim to adopt generative models to synthesize data for training segmentation models for such underrepresented modalities.
+Concretely, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within MRGen-DB, each pair comprises a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$).
+We leverage models trained on each dataset pair to generate target-domain samples for training segmentation models and assess our data engine from two aspects: (i) image generation quality and (ii) segmentation performance on the target-domain test set.
+Here, the test data for segmentation is strictly not used for training our generative data engines to prevent information leakage.
 Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [69] CycleGAN [70] UNSB [28] MRGen (Ours)
@@
 Baselines.
-For generation, we compare generated images from MRGen against three representative approaches: CycleGAN [69] and UNSB [27] for translating source-domain images to the target domain; and DualNorm [68] for exhaustive augmentation of source-domain images.
-For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data.
+For generation, we compare generated images from MRGen against three representative methods: CycleGAN [70] and UNSB [28] for translating source-domain images to the target domain; and DualNorm [69] for exhaustive augmentation of source-domain images.
+For segmentation, we assess models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data.
 We adopt nnUNet [23] and UMamba [41] as segmentation frameworks for all settings, except for DualNorm, which employs a customized UNet following their official codes.
-Comparisons with additional baselines are provided in Sec.
-D.2.
+More baseline results are provided in Sec.
+D.2 of the Appendix.
 Evaluation metrics.
-We employ distinct metrics for the evaluation: For image generation, we employ Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images.
-For segmentation models, we employ commonly used Dice Similarity Coefficient (DSC) [41] score to compare predicted masks with ground truth.
+For image generation, we leverage Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images.
+For segmentation models, we employ the commonly used Dice Similarity Coefficient (DSC) [42] score to compare predicted masks with ground truth.
 Considering segmentation consistency, we stack slices into 3D volumes, calculate the DSC for each organ individually, and average them as the final result.
@@
 For each mask, we generate 20 image candidates and select the best two satisfying the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively.
-Conditions for target-domain data synthesis are directly derived from region, organs information, and segmentation masks of source-domain data.
+Conditions for target-domain data synthesis are directly derived from the region, organs information, and segmentation masks of the source-domain data.
 Source Dataset Source Modality Target Dataset Target Modality DualNorm [69] UMamba [41] nnUNet [23]
@@
 Segmentation.
-As depicted in Table 3, we draw the following three observations: (i) significant discrepancies between different modalities challenge the generalization ability of nnUNet and UMamba models trained solely on source-domain data, leading to notably lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements in average DSC scores, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the highest DSC score in 8 out of 10 experiments, significantly outperforming others.
-Notably, MRGen consistently improves performance of nnUNet and UMamba, demonstrating the adaptability and robustness of its synthetic data across various segmentation architectures.
-More comparisons will be included in the Sec.
-D.2.
+As depicted in Table 3, we draw the following observations: (i) notable discrepancies among different modalities challenge the generalization ability of nnUNet and UMamba trained solely on source-domain data, leading to lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the best DSC score in 8 out of 10 experiments, significantly outperforming others.
+Notably, MRGen consistently improves the performance of nnUNet and UMamba, exhibiting the robustness and adaptability of its synthetic data across various segmentation architectures.
+More comparisons are included in Sec.
+D.2 of the Appendix.
 4.3 Qualitative Results
@@
 Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging.
-We visualize liver, right kidney, left kidney and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
+We visualize liver, right kidney, left kidney, and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
 Segmentation.
-As illustrated in Figure 5, despite significant appearance variations among distinct modalities, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data.
+As illustrated in Figure 5, despite significant appearance variations among distinct modalities, training on synthetic data from MRGen substantially improves segmentation accuracy across all organs.
 However, DualNorm, and segmentation models trained on data derived from CycleGAN and UNSB, yield unsatisfactory results.
@@
 4.4 Ablation Studies
-To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation task, as follows.
+To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation tasks, as follows.
 Generation.
@@
 As presented in Table 4, we can draw the following observations: (i) while SDM pretrained on natural images performs poorly on MRI, finetuning on MRGen-DB yields substantial improvement in both reconstruction and generation; (ii) MRGen, with higher latent dimensions (16 versus 4 of SDM) and the BiomedCLIP text encoder, achieves significantly better performance; and (iii) our templated text prompts further enable MRGen to better distinguish distinct modalities, regions, and organs, leading to superior synthesis quality.
-To this end, we employ high-capacity autoencoders trained on MRGen-DB, text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
+To this end, we employ high-capacity autoencoders trained on MRGen-DB, a text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
 Method AutoFilter $\mathcal{D}_{t}$ Image CHAOS-MRI [27] MSD-Prostate [3]
@@
 We train nnUNet with data generated by MRGen under various training and inference settings.
-As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without incorporating target-domain unannotated data during training, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain unannotated images and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
+As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without training on MR images from target domain, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain MR images (without annotations) and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
 5 Conclusion
@@
 We believe this will offer new insights into addressing the scarcity of medical data and annotations, holding clinical significance.
+Acknowledgments
+This work is supported by the National Key R & D Program of China (No. 2022ZD0161400).
 Appendix
@@
 $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$
-Here, $\mu_{\theta}$ represents the predicted mean of Gaussian distribution, derived from the estimated noise $\boldsymbol{\epsilon}_{\theta}$ as:
+Here, $\mu_{\theta}$ represents the predicted mean of the Gaussian distribution, derived from the estimated noise $\boldsymbol{\epsilon}_{\theta}$ as:
 $\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t))$
@@
 Appendix B Details of MRGen-DB & Synthetic Data
-In this section, we provide additional details about our collected and curated MRGen-DB dataset.
+This section provides additional details about our curated MRGen-DB dataset.
 In Sec.
 B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Sec.
-B.2, we present more comprehensive dataset statistics.
-Additionally, in Sec.
-B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training under each experimental setting.
+B.2, we present more comprehensive data statistics.
+Moreover, in Sec.
+B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training.
 B.1 Automatic Annotations
-We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and comprehensive clinically relevant information to distinguish distinct modalities, regions, and organs.
+We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and clinically relevant information to distinguish distinct modalities, regions, and organs.
 This process primarily consists of two precise and controllable components: human body region classification and modality explanation, which will be detailed as follows.
 Region classification.
-Considering the wide range and variability of abdominal imaging, we adopt the off-the-shelf BiomedCLIP [62] image encoder to encode all 2D slices and the BiomedCLIP text encoder to encode predefined text descriptions of six abdominal regions.
-Based on the cosine similarity between the image and text embeddings, the 2D slices are classified into these six categories, including Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region.
+Considering the wide range and variability of abdominal imaging, we adopt the off-the-shelf BiomedCLIP [63] image encoder to encode all 2D slices, and the BiomedCLIP text encoder to encode predefined text descriptions of six abdominal regions.
+Based on the cosine similarity between the image and text embeddings, the 2D slices are classified into one of the six categories, including Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region.
 For text encoding, we use a templated text prompt as input:
@@
 B.3 Synthetic Data Statistics
-This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings.
+This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings, as presented in Table 7.
 Concretely, we use mask annotations from the entire source-domain dataset (including both training and test sets) as input conditions to generate target-domain images, forming image-mask training pairs.
@@
 Otherwise, all low-quality generated samples are discarded to avoid noisy data.
-Finally, the statistics of target-domain training pairs generated by MRGen for each experimental setting are summarized in Table 7.
 Source Dataset Source Modality Target Dataset Target Modality # Slices ($\mathcal{D}_{s}$) # Synthetic Data
@@
 In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller.
-We do not explicitly adopt data balancing, as we empirically find that it does not lead to significant performance changes.
-For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [40].
+Our preliminary experiments show that MRGen remains robust to uneven data distribution; we therefore do not explicitly adopt data balancing in training.
+For segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [41].
 C.2 Autofilter Pipeline
@@
 Notably, these approaches, to varying degrees, rely on manually annotated target-domain segmentation masks or external datasets.
-Thus, they should be regarded as references only rather than fair comparisons with the aforementioned methods and our MRGen.
+Thus, they should be regarded as references only, rather than fair comparisons with the aforementioned methods and our MRGen.
 Few-shot methods.
@@
 D.2, we present quantitative comparisons with more baselines, further confirming the effectiveness and necessity of our proposed data engine.
+Then, in Sec.
+D.3, we present extra promising application prospects (cross-protocol generation) on paired CT and MRI datasets.
 Finally, in Sec.
-D.3, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
+D.4, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
 D.1 In-domain Generation
 Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domains.
-Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated data.
+Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real, manually-annotated data.
 This offers a feasible solution to address concerns about medical data privacy.
+Figure 7: Qualitative Results of In-domain Generation.
 Dataset Source Modality Target Modality $\mathcal{D}_{s}$ $\mathcal{D}_{t}$
@@
 We compare the performance of nnUNet [23] trained on real data versus synthetic data generated by MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
-Figure 7: Qualitative Results of In-domain Generation.
 Moreover, we provide visualizations of in-domain generation in Figure 7, qualitatively demonstrating that our MRGen can reliably perform controllable generation of a large number of samples within the training domain with mask annotations.
@@
 For computational efficiency, we primarily focus on comparing MRGen with some representative baselines, DualNorm [69], CycleGAN [70] and UNSB [28], across more dataset pairs in the main text for a comprehensive evaluation.
-D.3 More Qualitative Results
+Source Domain Target Domain DualNorm nnUNet UMamba
+$\mathcal{D}_{s}$ CycleGAN MRGen $\mathcal{D}_{s}$ CycleGAN MRGen
+AMOS22 (CT) CHAOS (T2) 19.78 0.11 6.75 22.50 0.05 8.06 26.73
+AMOS22 (CT) CHAOS (T1) 16.09 8.88 52.49 56.23 3.19 43.21 60.53
+MSD-Liver (CT) CHAOS (T2) 1.58 3.12 10.14 38.67 1.65 11.06 40.93
+Table 10: Quantitative Results (DSC score) on Cross-protocol settings (from CT to MRI).
+D.3 Extension to More Modalities
+Considering that evaluation on truly rare modalities is difficult due to limited ground truth annotations, we simulate such scenarios by restricting models from accessing target-modality labels in our experiments.
+Here, we also explore cross-modality synthesis (from CT to MRI) with AMOS22 [25], MSD-Liver [3], and CHAOS-MRI [27] datasets to further demonstrate MRGen’s potential for broader cross-protocol generation, as depicted in Table 10.
+D.4 More Qualitative Results
 In this section, we provide qualitative visualizations of more datasets, covering both image generation and segmentation.
 Image generation.
-As depicted in Figure 9, we present additional visualizations of controllable generation on target modalities lacking mask annotations.
-These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
+We present extra visualizations of controllable generation on target modalities lacking mask annotations in Figure 10, which demonstrate that our MRGen can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
 Image segmentation.
@@
 This validates that the samples generated by MRGen can effectively assist in training segmentation models, achieving impressive performance in previously unannotated scenarios.
+Figure 8: Failure Cases Analysis.
+Our proposed MRGen is not without limitations: (a) it may struggle to handle extremely small organ masks; (b) it occasionally produces false-negative samples, such as the unexpected synthesis of kidneys in the given example.
 Appendix E Limitations & Future Works
@@
 False-negative samples.
-Another challenge arises from the varying number of organs to be segmented on each slice.
-For instance, one slice may contain the liver, kidneys, and spleen, while another may include only the liver and spleen.
-This variability causes MRGen to occasionally generate additional segmentation targets not specified in the mask condition.
-For example, as illustrated in Figure 8 (b), kidneys are unexpectedly synthesized by MRGen, despite not being included in the mask conditions, leading to false negatives during the training of downstream segmentation networks.
-A feasible solution is to design a more comprehensive and robust data filtering pipeline to filter these false-negative samples.
-Alternatively, simple manual selection can serve as a quick and effective method to remove samples that do not meet the requirements.
-Figure 8: Failure Cases Analysis.
-Our proposed MRGen is not without limitations: (a) it may struggle to handle extremely small organ masks; (b) it occasionally produces false-negative samples, such as the unexpected synthesis of kidneys in the given example.
+Another challenge arises from the varying number of organs on each slice.
+For instance, one slice may contain liver, kidneys, and spleen, while another may include only liver and spleen.
+This variability causes MRGen to occasionally generate targets not specified in the mask condition.
+As depicted in Figure 8 (b), kidneys are unexpectedly synthesized, despite not being included in the mask conditions, leading to false negatives during the training of downstream segmentation networks.
+A feasible solution is to design a more robust data filtering pipeline to filter false-negative samples, and simple manual selection can also serve as a quick and effective method to remove the non-compliant samples.
+Figure 9: The validation of 3D extension feasibility of MRGen on $256\times 256\times 16$ volumes.
 E.2 Future Works
-Due to limited computational resources, we validate our proposed data engine on 2D slices in this paper, with trained segmentation models able to process 3D volumes slice-by-slice.
-However, we believe our idea can be seamlessly extended to 3D volume generation.
-With more computing in the future, we aim to develop a 3D version model for the community, further advancing cross-modality segmentation performance.
-Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
+Due to limited computational resources, we validate our data engine on 2D slices, with trained segmentation models able to process 3D volumes slice-by-slice.
+However, our idea can be seamlessly extended to 3D volume generation with more computing in the future to further advance cross-modality segmentation performance.
+Here, we provide a preliminary validation on $256\times 256\times 16$ volumes, as depicted in Figure 9.
+While the results are not fully optimized due to limited computations, they already demonstrate promising inter-slice consistency, indicating the feasibility of extending MRGen to 3D synthesis.
+Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
 Figure 10: More Qualitative Results of Controllable Generation.
-We present images from source domains $\mathcal{D}_{s}$ and target domains $\mathcal{D}_{t}$ for reference.
+We present images from source domains ($\mathcal{D}_{s}$) and target domains ($\mathcal{D}_{t}$) for reference.
 Here, specific organs are contoured with colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver in LiQA and CHAOS-MRI datasets.